### v1

# Conceptwm: A Diffusion Model Watermark for Concept Protection

Abstract
The personalization techniques of diffusion models succeed in generating specific concepts but also pose threats to copyright protection and illegal use. Model Watermarking is an effective method to prevent the unauthorized use of subject-driven or style-driven image generation, safeguarding concept copyrights. However, under the goal of concept-oriented protection, current watermarking schemes typically add watermarks to all images rather than applying them in a refined manner targeted at specific concepts. Additionally, the personalization techniques of diffusion models can easily remove watermarks. Existing watermarking methods struggle to achieve fine-grained watermark embedding with a few images of specific concept and prevent removal of watermarks through personalized fine-tuning. Therefore, we introduce a novel concept-oriented watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models. We conduct extensive experiments and ablation studies to verify our framework. Our code is available at https://anonymous.4open.science/r/Conceptwm-4EB3/.
1 Introduction
Recent advancements in diffusion models have revolutionized the field of image generation, making it possible to create realistic images from simple text description inputs. With the growing demand for generating highly personalized and customized content, the personalization techniques [29, 11, 10, 18] of diffusion models have achieved significant progress in generating novel concept descriptions (e.g., unique visual themes, motifs, stylistic preferences, or characters) with a few images of specific themes or styles. Through powerful personalization tools like DreamBooth [27] and Lora [14], anyone, even without design skills, can create images of specific concepts using just a few personal images and simple text inputs.
Figure 1: (a) Post-processing Watermark. (b) U-net Finetuning Watermark. (c) VAE-Finetuning Watermark. (d) Latent-based Watermark. (e) Ours.
However, the use of personalization techniques introduces significant risks, particularly the potential for concept misuse or leakage. Untrustworthy users can exploit these tools to create unauthorized concepts related to celebrities or artworks, leading to deepfake incidents [15], the spread of fake news and online spam [4], and serious copyright violations, as the original creators’ intellectual property and creativity are undermined.
To address the ownership and accountability issues of specific concepts in generative models, recent research has introduced the task of concept watermarking [7], which aims to embed watermark information to the concept and extract the watermark from the subsequent generated images based on the watermarked concept. In practice, the model owner typically possesses a pre-trained diffusion model and seeks to enable the generation of images for a new concept. The concept owner provides images with a specific concept to fine-tune the existing diffusion model while embedding watermark information related to the concept, enabling the model to generate images containing the new concept and allowing the watermark decoder to extract watermark information from the generated images.
Even though adding watermarks to generated images has brought increasing attention [17, 36, 41, 37, 9, 43, 34, 8, 6], it is non-trivial for existing watermarking solutions to achieving concept watermarking. On one hand, most existing methods [34, 9, 43] typically watermark the whole images rather than images of specific concepts, so that they are inapplicable in the concept watermarking scenario. Fine-grained watermarking mechanisms for specific concepts has rarely addressed by prior work. The major challenge is that imgae watermarking concept watermarks must not only exist within a specific image, but also remain present across various image variations generated by diffusion model. Dissimilar to traditional image watermarks embedded at the pixel level, the concept watermarks appears a higher-level technical barrier as it associates the watermarks with high-levels of sematic features related to specific concepts. Since the amount of images with a specific concept generally is limited for personalization model fine-tuning by injecting a new concept, it is a great challenge for existing methods which embed watermarks with a large volume of training samples.
On the other hand, robustness is another challenge for existing watermarking methods, when considering diverse adversaries in the concept protection scenarios. Recent work also showed that existing methods failed in defending model purification attacks [45, 21], as adversaries can bypass watermarking schemes by generating various watermark-free images through personalized fine-tuning. Balancing adversariality, image quality, and watermark fidelity becomes a technical obstacle. One of alternatives for increasing watermark adversariality is to inject imperceptible protective adversarial perturbations into the protected images, but this idea generally has a negative impact on both image quality and watermark fidelity [19, 28, 31]. Therefore, finding out a watermarking solution to ensuring concept tracing while addressing threats from adversaries has an urgent demand.
There are three core properties needed for satisfying the requirement of concept watermarking, namely, refined nature, robustness, and adversarial properties. Refined nature: The knowledge of a specific concept is learned by the denoising network (U-net), which is the core component of the diffusion model. Adding a watermark to a specific concept should be done with the least amount of training data and training steps possible, while ensuring that the fidelity before and after adding the watermark remains unaffected. Robustness: The watermark should remain stable across various image processing operations and under different generation parameters. Adversarial properties: The generated concept-specific images should not be subject to “jailbreaking” via personalized fine-tuning.
To address the challenges outlined above, we propose a novel concept watermarking framework ConceptWm. We consider that a refined watermark training and Fidelity Preserving Perturbation are two critical factors in the context of concept watermarking. To achieve the concept watermark embedding, we propose Concept-oriented Watermarking Finetuning, which ensures the rapid learning of watermark embedding patterns in diffusion models with limited data and training budget. To ensure the robustness of the watermark against various image processing attacks and facilitate the learning of the watermark diffusion model, our Watermark Components Pretraining Module incorporates the watermark into the latent space and employs a distortion layer to ensure the watermark can withstand various image processing attacks. To effectively balance adversarial perturbation, watermark fidelity, and image quality, while mitigating the risk of adversarial attacks bypassing the watermark through personalized fine-tuning, we introduce the Fidelity Preserving Perturbation Modulation. This approach employs an alternating optimization strategy, iteratively refining both the model parameters and adversarial perturbations, thus attaining an optimal equilibrium between adversarial robustness, image quality, and watermark fidelity.
Our contributions are as follows: (1) We have proposed a novel concept-driven watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models. We consider that a refined watermark training and fidelity preserving perturbation are two critical factors in the context of concept watermarking, which offers a fresh perspective on how diffusion model watermarking can be applied in real-world scenarios. (2) We propose a mechanism named Fidelity Preserving Perturbation Modulation for endowing concept watermarks with adversarial properties. Through the alternating optimization of model training and adversarial algorithms, it effectively balances adversarial robustness, image quality, and watermark fidelity. (3) Experiments demonstrate that Conceptwm effectively integrates adversarial perturbations into watermarks and maintains better robustness and accuracy in concept watermark embedding and extraction compared with existing 4 baselines over 6.3% image processing attacks.
2 Related Work
Personalization Techniques for Diffusion Models. Given the substantial computational requirements of training Stable Diffusion from scratch, many methods aim to inject specific concepts into the model through fine-tuning. Currently used fine-tuning techniques include Textual Inversion [10], DreamBooth [27] and LoRA (Low-Rank Adaptation)[14]. Textual Inversion focuses solely on training text embeddings during fine-tuning, allowing concepts to be injected into the text encoder without altering the weights of the U-Net. DreamBooth, on the other hand, fine-tunes the entire U-Net portion of the Stable Diffusion model. Unlike conventional text-to-image fine-tuning, DreamBooth employs prior loss during training to prevent overfitting. LoRA achieves quick and lightweight fine-tuning of Stable Diffusion by training incremental weights in the attention layers of the U-Net.
Model Watermarking for Latent Diffusion Models. WDM [43] trains an autoencoder to stamp a watermark on all training data before re-training the generator from scratch. Aqualora [8] propose a watermark Lora method to watermark all the images. VAE-based methods [36, 17, 9] fine-tune VAE-Decoder to ensure that all generated images contain the watermark. However, the aforementioned watermarking schemes apply watermarks to all generated images, leaving a gap in the protection of specific styles. Meanwhile, A large number of images is often required for the model to learn the watermark embedding pattern, but concept-oriented images for finetuning are often limited.
Protective Perturbation against Stable Diffusion. To safeguard personal images, such as faces and artwork, from potential infringement during the fine-tuning of Stable Diffusion, recent studies focus on disrupting the fine-tuning process by adding imperceptible protective noise to these images. AdvDM [19] introduces a direct adversarial attack on Stable Diffusion by maximizing the Mean Squared Error loss during optimization. This method employs adversarial noise to safeguard personal images. Anti-DreamBooth [31] integrates the DreamBooth fine-tuning process of Stable Diffusion into its framework, designing a bi-level min-max optimization approach to generate protective perturbations. Additionally, other studies [35, 38, 44, 46] have investigated the use of similar adversarial perturbation techniques to generate protective noise for images.
3 Preliminaries
Prompt-based Latent Diffusion Model. Conditional denoising model, denoted as $\epsilon_{\theta}(x,t,c)$, can generate images based on a given text condition $c$. $x_{t}$ represents the latent representation at a particular timestep t within the set $\{1,2...,T\}$. During training, a loss function $\mathcal{L}_{cond}$ is used to guide the diffusion model in denoising the latent representations, $x_{t}=\sqrt{\bar{\alpha}_{t+1}}x_{0}+\sqrt{1-\bar{\alpha}_{t+1}}\epsilon_{\theta}(x_{t})$. Given the input image $x_{0}$, each variable $x_{t}$ is constructed via injecting noise at corresponding timestep $t$. The objective can be expressed as follows:
$\mathcal{L}_{cond}(\theta,x_{0})=\mathbb{E}_{x_{0},t,c,\epsilon\in\mathcal{N}(0,1)}\|\epsilon-\epsilon_{\theta}(x_{t+1},t,c)\|_{2}^{2}.$
DreamBooth. DreamBooth employs the base loss of diffusion models in Eq.1. It also incorporates a prior preservation loss to avoid overfitting and text-shifting issues, particularly when only a limited set of instance examples is available. More precisely, it uses a generic prior prompt such as “a photo of [class]” and learn the concept with the prompt such as “a photo of <concept >[class]”. The objective can be expressed as follows:
$\mathcal{L}_{ft}(\theta,x_{0})=\mathbb{E}_{x_{0},t}\|\epsilon-\epsilon_{\theta}(x_{t+1},t,c)\|_{2}^{2}+\lambda\|\epsilon^{\prime}-\epsilon_{\theta}(x_{t+1}^{\prime},t,c_{cls})\|_{2}^{2},$
where $\epsilon$ and $\epsilon^{\prime}$ are both sampled from the standard normal distribution $\mathcal{N}(0,\mathbf{I})$, the noisy variable $x_{t+1}^{\prime}$ is generated from the original stable diffusion model $\theta_{\text{ori}}$ using the prior prompt $c_{\text{cls}}$. The term $\lambda$ emphasizes the importance of the prior term in the generation process.
Adversarial attacks. The goal of adversarial attacks is to find imperceptible perturbations to the input image that mislead the behavior of a given model. Typical research usually focuses on classification tasks, where, for a model $f$, an adversarial example $x^{\prime}$ of an input image $x$ is generated such that it is visually indistinguishable, while causing a misclassification $y=f(x^{\prime})$ The perturbation is typically constrained to lie within an $\eta$-ball to ensure minimal visual difference, specifically satisfying $\|x^{\prime}-x\|_{p}<\eta$. Projected Gradient Descent (PGD) [23] is a widely used attack method that relies on an iterative optimization process. The procedure to update and predict $\tilde{x}$ for an untargeted attack is as follows:
$\tilde{x}^{t+1}=\Pi_{(x,\eta)}\left(\tilde{x}^{t}+\alpha\cdot sign(\nabla_{x}\mathcal{L}(f(x),y))\right),$ (3)
where $\Pi_{x,\eta}(z)$ restricts the pixel values of $z$ within an $\eta$-ball around the original values in $x$.
4 Methodology
Figure 2: Framework of the proposed Conceptwm watermarking scheme for Latent Diffusion Model.
4.1 Motivation
Challenges of existing methods: (1) How to achieve concept-oriented watermark fine-tuning of U-net with limited data and training budget? We will solve this via the framework design in Sec. 4.4. (2) Existing watermarking methods lack protective mechanisms to prevent watermark ”jailbreaking” through personalized fine-tuning. Integrating the existing protective perturbation method into the watermarking mechanism and embedding it into the model while ensuring good generation quality is challenging. This will be solved in Sec. 4.5.
4.2 Overview
The main objective of our method is to protect images of the specific concept generated by the diffusion model by embedding adversarial noise and the watermark pattern into the U-net component of the diffusion model. Figure 2 illustrates an overview of our approach. The Watermark Components Pretraining module implements watermarking in the latent space and enhances the robustness of the watermark through a distortion layer. Fidelity Preserving Perturbation Modulation implements adversarial perturbations to the image while balancing watermarking, image quality, and adversarial properties. Concept-oriented Watermark Fine-tuning enables learning for specific concepts and integrates adversarial watermarks through fine-tuning.
4.3 Watermark Components Pretraining
Existing image watermarking methods often face significant differences between the watermark embedding pattern and the latent space, which causes the watermark information to be disrupted or even lost when transformed into the latent space by the VAE encoder or diffusion regeneration. This issue has been discussed in related watermarking attack methods [1, 3, 5]. Our objective is to embed a robust watermark within the image that is both resilient and conducive to efficient learning by the U-Net architecture.
The U-Net learns the distribution and noise of the image in the latent space, and embedding the watermark into the latent space enables the model to learn the consistency between the watermark and the generated images. In this step, the watermarked images are processed through the message decoder to evaluate the discrepancy between the decoded watermark information and the original content of the watermarked image. The message loss is the Binary Cross Entropy (BCE) between the message $m$ and the sigmoid $\sigma(m^{\prime})$.
$\mathcal{L}_{m}=-\sum_{k=0}^{n-1}m_{k}\log\sigma(m_{k}^{\prime})+(1-m_{k})\log(1-\sigma(m_{k}^{\prime}))$ (4)
$I_{0}$ is the original image. $I_{r}$ is the reconstructed image. $m$ is secret message (i.e.,watermark). To ensure visual consistency, we use the LPIPS loss function [40] to evaluate the global loss between the watermarked image and the reconstructed image. To avoid excessive modification of local image regions, we select the most prominent parts of the image and calculate channel loss.
$R$ $es(\mathbf{I}_{0},\mathbf{I}_{r})=\frac{1}{3}\sum_{c=1}^{3}|{\mathbf{I}_{c}}^{0}(i,j)-{\mathbf{I}_{c}}^{r}(i,j)|$ (5)
$\mathcal{L}_{channel}=\sum_{k=1}^{K}\left[\operatorname{Top-k}\left(\operatorname{Res}(\mathbf{I}_{0},\mathbf{I}_{r})\right)\right]$
Our training objective can be summarized as follows, where $\lambda$ and $\mu$ represent the coefficients:
$\mathcal{L}_{total}=\mathcal{L}_{m}+\lambda\mathcal{L}_{pips}+\mu\mathcal{L}_{channel}$ (6)
The trained message-encoder and message-decoder will collaborate with the Fidelity Preserving Perturbation Modulation module to assist the U-Net in learning the adversarial watermark.
4.4 Concept-oriented Watermark Fine-tuning
At this stage, the watermark is embedded into specific concepts and then injected into the U-Net. The main challenges are as follows: The fine-tuned images for specific concepts are often limited, making it difficult for the model to learn the specific watermark patterns; Due to the introduction of adversarial perturbations, directly using the DreamBooth training paradigm will be affected by protective perturbations, preventing the model from learning the concept.
The limited number of specific concept samples and adversarial perturbation often result in failures when directly adding an adversarial watermark into images and deploying DreamBooth. This can prevent the watermark embedding pattern from being learned effectively and lead to quality degradation. Our core idea is to decouple concept learning from adversarial watermark learning. In the DreamBooth stage, we learn the specific concept, and by adding adversarial noise to the same image and also using the image without the adversarial noise, the model learns to differentiate between them. The training objective of this stage is as follows:
$\mathcal{L}_{\mathrm{ft}}(\theta)$ $:=\mathbb{E}_{t,c,x}\Big{[}\|\boldsymbol{\epsilon}_{\theta}(x_{t}+\varphi,t,c)-\boldsymbol{\epsilon}_{ori}(x_{t},t,c\big{)}\|^{2}\Big{]}$ (7)
$\varphi$ represents the adversarial perturbation containing the watermark generated through Fidelity Preserving Perturbation Modulation, $\epsilon_{ori}$ denotes the fixed model, and $\epsilon_{\theta}$ refers to the trained model.
4.5 Fidelity Preserving Perturbation Modulation
The goal of this fine-tuning is to maximize the loss of $L_{cond}$ in Eq. 1 and minimize the loss of $L_{ft}$ 2. The adversarial requirement for watermarking is to disrupt the learning process of DreamBooth. Since the DreamBooth model tends to overfit on adversarial images, this can be exploited to make it perform worse when reconstructing clean images.
$\sigma^{*(i)}$ $=\underset{\delta^{(i)}}{\arg\max}\mathcal{L}_{\text{cond }}(\theta^{*},x^{(i)}),\forall i\in\{1,..,N_{ft}\},$ (8)
[…]
and $\|\sigma^{(i)}\|_{j}\leq\eta\quad\forall i\in\{1,..,N_{ft}\}$
where $x^{i}$ denotes the clean samples. Meanwhile, we need to ensure that the image quality and watermark loss are as minimal as possible. Our solution is to use watermark-free images for one-step DreamBooth to reduce the DreamBooth loss $\mathcal{L}_{\text{cond}}$ 1, while simultaneously applying PGD to enhance the adversarial loss $L_{ft}$. 2.
$\theta$ $\leftarrow\arg\min_{\theta}\sum_{x\in\mathcal{X}_{A}}\mathcal{L}_{ft}(\theta,x^{i})$ (9)
$\theta^{\prime}$ $\leftarrow\theta.\text{clone}()$
$\varphi^{(i)}$ $\leftarrow\arg\max_{\delta}\mathcal{L}_{cond}(\theta^{\prime},x^{(i)}+\sigma_{1}+\sigma_{2}^{(i)})$
Considering that the adversarial perturbation will affect the watermark, we also incorporate the decoded message loss as a constraint to optimize the perturbation.
$\mathcal{L}_{msg}=BCE(decoder(x+\sigma_{1}+\sigma_{2}^{(i)}),m_{target})$ (10)
where $\sigma_{1}$ denotes the residual of the watermarked image and original image. The Fidelity Preserving Perturbation Modulation outputs an adversarial watermark $\varphi$ as the parameter for Concept-oriented Watermark Fine-tuning to facilitate learning.
Figure 3: Demonstration of the generated images. Below each image, we display the corresponding Bit Accuracy, along with the image quality metrics PIQE and NIQE.
Method Concept Oriented Concept Protection Low Data Requirement Fidelity Robustness
FID CLIP BitAcc BitAcc(Adv.) TPR TPR (Adv.)
None - - - 24.37 0.367 - - - -
Post-Processing
DwtDctSvd $\times$ $\times$ 25.01 0.363 100.0 71.47 1.000 0.367
RivaGan $\times$ $\times$ 24.72 0.363 98.47 83.21 0.994 0.651
Latent-based
Tree-Ring $\times$ $\times$ 25.69 0.364 - - 1.000 0.821
Merging Component
StableSig. $\times$ $\times$ $\times$ 24.92 0.362 98.49 81.94 0.993 0.848
Ours 24.58 0.365 95.68 89.32 0.994 0.884
Table 1: A comparison between our method and previous watermarking techniques is presented. The capacities for DWT-DCT-SVD, RivaGAN, StableSignature, and our method are 48 bits, 32 bits, 48 bits, and 48 bits, respectively. We control the false positive rate (FPR) at $10^{-6}$ and evaluate the true positive rate (TPR). Since Tree-ring is a zero-bit watermark, bit accuracy cannot be calculated in this case. ”Adv.” (Adversarial) refers to the average performance of images under various distortions.
5 Experiment
METHODS DISTORTIONS
Contrast Brightness BLUR Gauss Noise JPEG Crop&Rescale Vae.C Denoising Average
Dwtdctsvd 88.41 86.80 92.42 65.87 83.32 54.02 55.74 49.53 72.01
RivaGan 97.96 94.67 94.89 91.09 90.68 94.17 70.62 58.64 86.59
Stablesig. 97.86 95.48 89.88 78.94 84.40 97.47 69.94 52.58 83.31
Ours 94.26 93.46 93.68 92.42 93.63 90.26 90.12 84.63 91.34
Dwtdctsvd 0.894 0.752 0.943 0.054 0.734 0.008 0.010 0.000 0.424
RivaGan 0.989 0.956 0.979 0.926 0.928 0.956 0.284 0.010 0.754
Stablesig. 0.987 0.985 0.905 0.369 0.821 0.989 0.293 0.000 0.669
Tree-ring 1.000 1.000 0.954 0.701 0.905 0.121 0.958 1.000 0.830
Ours 0.968 0.951 0.947 0.958 0.949 0.912 0.910 0.734 0.916
Table 2: Our method’s performance against various types of attacks is compared with other watermarking schemes. The attacks include (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression [5] and (8) Diffusion-based Reconstructive Attack [42].
Settings BitAcc(%) TPR
Steps 25 0.917 0.921
50 0.936 0.946
100 0.936 0.951
Sampler DDIM 0.929 0.939
DPM-M 0.938 0.972
EULER 0.922 0.941
HEUN 0.947 0.962
Guidance Scale 5 0.932 0.948
7.5 0.917 0.932
10 0.933 0.947
Generation Size 512 0.933 0.951
768 0.812 0.748
VAE V2.1 0.939 0.954
V1.4 0.901 0.913
Table 3: The table presents the results of our method under different experimental parameters. Our concept watermarking approach remains effective across various generation steps, different samplers, VAE versions, and generated shapes.
TRAIN TEST “a photo of <concept >person” “a portrait of <concept >person”
PIQE↓ NIQE↓ CLIP↑ FID↓ PIQE↓ NIQE↓ CLIP↑ FID↓
MODEL MISMATCH v2.1 v1.4 63.18 6.87 0.287 87.21 54.7 6.79 0.272 84.53
v1.4 v2.1 59.82 6.72 0.293 72.18 57.8 6.37 0.292 66.28
LORA v2.1 v2.1 53.82 5.72 0.306 69.48 55.82 5.37 0.311 70.98
Textual Inversion v2.1 v2.1 36.87 4.96 0.294 66.47 39.58 4.48 0.285 56.83
Table 4: An experiment on generating images with adversarial watermarks and fine-tuning them with different model versions and methods for personalization. The effectiveness of our adversarial watermark defense through personalized fine-tuning is observed across different model versions and personalized methods.
PSNR SSIM
BitAcc
BitAcc
GaussianNoise 33.26 0.928 0.962 0.934
ColorJitter 34.78 0.939 0.979 0.972
GaussianBlur 33.69 0.931 0.967 0.948
Jpeg 32.26 0.933 0.969 0.882
CROP&SCALE 30.58 0.917 0.959 0.593
Random Mask 32.19 0.915 0.969 0.908
Combination 29.08 0.889 0.954 0.528
Table 5: The impact of various attacks on image quality in the distortion layer in the watermark components pretraining.
5.1 Implementation Details
Datasets. To evaluate the effectiveness of the proposed methods. For subject-driven generation, we select the face dataset CelebA-HQ [47]. For style-driven generation, we select 12 artists with distinct styles from WikiArt  [^1] https://www.wikiart.org, with each artist contributing 10 artworks that are consistent in terms of both artistic style and time period. Each image is resized and center-cropped to a resolution of 512 × 512.
Training Configurations. By default, we use the Stable Diffusion (v2.1) as the pretrained generator. Unless stated otherwise, the training instance prompt for subject-driven generation is ”a photo of <concept >person” and the prior prompt is ”a photo of person”. The training instance prompt for style-driven generation is ”a painting in the style of <concept >” and the prior prompt is ”a painting”. For PGD, we use $\alpha$ = 1e-5 and the default noise budget $\eta$ = 1e-3. The learning rate of the dreambooth is set 5e-6.
Evaluation Metrics. We evaluate the detection performance using the true positive rate (TPR) at a false positive rate (FPR) of 10e-5. For traceability performance, we assess bit accuracy. To measure the quality of image generation, we compute the Peak Signal-to-Noise Ratio (PSNR) [13] and Structural Similarity Index (SSIM) [33] to evaluate image distortion; and the Natural Image Quality Evaluator (NIQE) [24] and Perceptual Image Quality Evaluator (PIQE) [32] for evaluating overall image quality.
Watermarking baselines. We select four typical baselines: two official watermark of Stable Diffusion [26] called DwtDctSvd [6], and RivaGAN [39],Vae-based method Stable Signature [9] and Latent-based method Tree-ring [34].
5.2 Watermark Performance Comparison
We evaluate our method against six typical baselines on two tasks: (1) Detection: All methods are tested as single-bit watermarks with a unified watermark, where the False Positive Rate (FPR) is set to 1%, and we assess the True Positive Rate (TPR) on 1,000 watermarked images. (2) Traceability: Except for the single-bit watermark Tree-Ring, all other methods are tested as multi-bit watermarks. In this experiment, we simulate a scenario with 1,000 model users, each requiring a distinct watermark for model tracing. Each user generates 10 images, creating a dataset of 10,000 watermarked images. During testing, we calculate the Bit Accuracy by comparing the watermark of each user with the watermarked images. The user with the highest Bit Accuracy is considered the traced user. The comparison results, shown in Table 1, demonstrate that our watermarking method achieves superior robustness, significantly outperforming the baselines in both tasks. In terms of Bit Accuracy, our method outperforms the best baseline by around 6.3%. This improvement is due to the widespread diffusion of the watermark throughout the latent space, creating a strong binding between the watermark and the image semantics.
5.3 Fidelity
Table 1 presents a comparison of our method with other baseline methods. For fidelity evaluation metrics, we use the Fréchet Inception Distance (FID) [12] for assessing image diversity and semantic relevance calculated on 5000 images from the COCO 2017 validation set [20] to assess image quality. Additionally, we also use CLIP score [25], a method for measuring image similarity, which provides results more aligned with human judgment.
Main Results. The performance of this method is relatively lower compared to other methods. It outperforms other methods in metrics like CLIP and FID, with minimal performance difference compared to the original model (FID difference of 0.21, CLIP difference of 0.02). This method is a localized fine-tuning of the model and does not significantly impact the model’s overall performance. Due to the smaller dataset and the more refined watermarking method, our method has a smaller impact on model generation quality compared to other methods.
5.4 Watermark Robustness against Attack
Image Processing. We select eight representative types of image-level noise to evaluate the robustness of image processing. The main results are shown in Table 2.
Main Results. For each image processing attack, we report the average bit accuracy in Table 2. We see that our Conceptwm is indeed robust across all the transformations with the bit accuracy all above 0.9. Conceptwm remarkably outperforms existing multi-bit watermarks on the average performance with more than 6.3% boost compared with the state-of-the-art (SoTA) results. Since our method embeds the watermark into the entire latent space and trains the distortion layer, it has a clear advantage over contrastive watermarking schemes when facing various image processing attacks. Please refer to the Supplementary Materials for more results.
5.5 Concept Protection Evaluation
Model Mismatching. We provide an example of transferring adversarial watermark trained on SD v1.4 and SD v2.1 to defend DreamBooth models trained from v2.1 and v1.4 in Table 4. The results show that our method still works effectively across different versions of the Diffusion Model.
Textual Inversion and DreamBooth with LoRA. Textual Inversion teaches new concepts by adjusting a word vector, rather than fine-tuning the entire model. LoRA, on the other hand, employs low-rank weight updates to enhance memory efficiency and is widely adopted within the community. The relevant results are shown in the Table 4. Our method shows better performance in mitigating the degradation effect on generation with LoRA. Since Textual Inversion does not inject concepts directly into the model, the reduction in quality is not significant and the semantic quality of the generation decrease.
Main Results. The results in Table 4 demonstrate that adversarial properties of our watermark is preserved across different versions of Stable Diffusion. When the model is mismatched, PIQE exceeds 57.8 and NIQE exceeds 6.37, indicating that our adversarial properties are inheritable across different versions of the diffusion model. Our method ensures that images generated with LoRA have a PIQE higher than 53.82, while the semantic quality metrics of images generated with Textual Inversion, measured by the CLIP score and FID, are 66.47 and 0.292. It effectively resists other types of personalization fine-tuning attacks, such as LoRA and Textual Inversion.
5.6 Ablation Studies.
Prompt Mismatching. Considering that users generate relevant images based on various prompts combined with specific concepts, we generate images for specific concepts under different prompt conditions. The main results are shown in Table 3, which shows that our watermarking method achieves higher bit extraction accuracy for images similar to the training set. When the image deviates significantly from the set instance prompt, our watermarking method still maintains good precision. Furthermore, the generated watermark exhibits a certain level of adversarial robustness, as the quality of images trained under this watermark is reduced, which is beneficial for the protection of specific concepts in diffusion models.
Distortion Layer. We evaluated the impact of various attacks on image quality in the Distortion Layer in the watermark components pretraining. The main results are in Table 5, which indicates that cropping and the mask adversarial layers have a significant impact on image quality, as they require a stronger watermark to be stably embedded within a smaller space.
Inference Steps. Users may influence watermark extraction under different inference parameters. We generated images of specific concepts at different inference steps and performed watermark extraction. The results, shown in Table 3, indicate that the inference steps have little impact on watermark extraction.
Guidance Scales. Larger guidance scales result in more faithful of the generated image adherence to prompts. We evaluate the impace of guidance scale in the Table 3. The results show that our method maintains good stability across various values of the Guidance scale.
Sampling Method. We evaluate our watermark on DDIM [30], DPM-Solver multi-step [22], Euler and Heun samplers [16]. The results show that our watermark is minimally affected by the choice of sampler.
Different VAE Decoder. An adversary may interfere with watermark decoding by replacing the VAE decoder. To evaluate the impact of the generated images on the watermark, we used different types of VAEs including stable diffusion V1.4 and V2.1 for decoding. The results in Table 3 indicate that changing the VAE does not have a significant impact on the watermark extraction, which is a clear advantage over VAE-based schemes, as the watermark cannot be preserved when the VAE is replaced in VAE-based methods.
6 Discussion
Limitations. First, our method exhibits limited detection performance under certain prompts. As shown in Figure 3, when the model generates poor results for specific concepts under certain prompts, the detection accuracy cannot be guaranteed. This issue requires more precise and improved personality techniques or larger datasets and prompts. Second, the adversarial perturbations generated by our proposed method have limited effectiveness against the latest purification methods [45, 2]. Since our approach involves embedding concepts, watermarks, and adversarial perturbations into the diffusion model, there are more constraints to consider, which makes our adversarial perturbations more limited compared to the state-of-the-art image adversarial perturbation methods [19, 31]. In the future, we will explore ways to better embed adversarial perturbations into the Diffusion model.
7 Conclusion
This paper reveals the potential threats posed by diffusion model personalization techniques and proposes an adversarial watermarking method to protect specific concepts. Our solution integrates adversarial perturbations and watermarks into the diffusion model, ensuring that images generated by the diffusion model for specific concepts are traceable, and any personalized model trained on these images will produce low-quality results. We have evaluated the stability and adversarial effectiveness of the watermarking method we designed, demonstrating that our defense remains effective under adverse conditions. In the future, our goal is to improve the quality and robustness of the watermark and consider more potential threats.

### v2

@@
-Conceptwm: A Diffusion Model Watermark for Concept Protection
-Abstract
+ConceptWM: Watermarking Diffusion Models for Visual Concept Copyright Protection
+Abstract.
 The personalization techniques of diffusion models succeed in generating specific concepts but also pose threats to copyright protection and illegal use.
@@
 Therefore, we introduce a novel concept-oriented watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models.
+We introduce Fidelity-preserving Latent Watermarking (FLW) to generate latent watermarks based on image characteristics and the Adversarial Watermarking Modulation module to prevent ”jailbreaking” via personalized finetuning.
+To enhance U-Net’s efficiency in learning watermark patterns with limited samples, we propose Efficient Concept Watermark Finetuning, which alternates optimization of model parameters for both watermark embedding and concept learning.
 We conduct extensive experiments and ablation studies to verify our framework.
 Our code is available at https://anonymous.4open.science/r/Conceptwm-4EB3/.
-1 Introduction
-Recent advancements in diffusion models have revolutionized the field of image generation, making it possible to create realistic images from simple text description inputs.
-With the growing demand for generating highly personalized and customized content, the personalization techniques [29, 11, 10, 18] of diffusion models have achieved significant progress in generating novel concept descriptions (e.g., unique visual themes, motifs, stylistic preferences, or characters) with a few images of specific themes or styles.
-Through powerful personalization tools like DreamBooth [27] and Lora [14], anyone, even without design skills, can create images of specific concepts using just a few personal images and simple text inputs.
-Figure 1: (a) Post-processing Watermark. (b) U-net Finetuning Watermark. (c) VAE-Finetuning Watermark. (d) Latent-based Watermark. (e) Ours.
-However, the use of personalization techniques introduces significant risks, particularly the potential for concept misuse or leakage.
-Untrustworthy users can exploit these tools to create unauthorized concepts related to celebrities or artworks, leading to deepfake incidents [15], the spread of fake news and online spam [4], and serious copyright violations, as the original creators’ intellectual property and creativity are undermined.
-To address the ownership and accountability issues of specific concepts in generative models, recent research has introduced the task of concept watermarking [7], which aims to embed watermark information to the concept and extract the watermark from the subsequent generated images based on the watermarked concept.
-In practice, the model owner typically possesses a pre-trained diffusion model and seeks to enable the generation of images for a new concept.
-The concept owner provides images with a specific concept to fine-tune the existing diffusion model while embedding watermark information related to the concept, enabling the model to generate images containing the new concept and allowing the watermark decoder to extract watermark information from the generated images.
-Even though adding watermarks to generated images has brought increasing attention [17, 36, 41, 37, 9, 43, 34, 8, 6], it is non-trivial for existing watermarking solutions to achieving concept watermarking.
-On one hand, most existing methods [34, 9, 43] typically watermark the whole images rather than images of specific concepts, so that they are inapplicable in the concept watermarking scenario.
-Fine-grained watermarking mechanisms for specific concepts has rarely addressed by prior work.
-The major challenge is that imgae watermarking concept watermarks must not only exist within a specific image, but also remain present across various image variations generated by diffusion model.
-Dissimilar to traditional image watermarks embedded at the pixel level, the concept watermarks appears a higher-level technical barrier as it associates the watermarks with high-levels of sematic features related to specific concepts.
-Since the amount of images with a specific concept generally is limited for personalization model fine-tuning by injecting a new concept, it is a great challenge for existing methods which embed watermarks with a large volume of training samples.
-On the other hand, robustness is another challenge for existing watermarking methods, when considering diverse adversaries in the concept protection scenarios.
-Recent work also showed that existing methods failed in defending model purification attacks [45, 21], as adversaries can bypass watermarking schemes by generating various watermark-free images through personalized fine-tuning.
-Balancing adversariality, image quality, and watermark fidelity becomes a technical obstacle.
-One of alternatives for increasing watermark adversariality is to inject imperceptible protective adversarial perturbations into the protected images, but this idea generally has a negative impact on both image quality and watermark fidelity [19, 28, 31].
-Therefore, finding out a watermarking solution to ensuring concept tracing while addressing threats from adversaries has an urgent demand.
-There are three core properties needed for satisfying the requirement of concept watermarking, namely, refined nature, robustness, and adversarial properties.
-Refined nature: The knowledge of a specific concept is learned by the denoising network (U-net), which is the core component of the diffusion model.
-Adding a watermark to a specific concept should be done with the least amount of training data and training steps possible, while ensuring that the fidelity before and after adding the watermark remains unaffected.
-Robustness: The watermark should remain stable across various image processing operations and under different generation parameters.
-Adversarial properties: The generated concept-specific images should not be subject to “jailbreaking” via personalized fine-tuning.
-To address the challenges outlined above, we propose a novel concept watermarking framework ConceptWm.
-We consider that a refined watermark training and Fidelity Preserving Perturbation are two critical factors in the context of concept watermarking.
-To achieve the concept watermark embedding, we propose Concept-oriented Watermarking Finetuning, which ensures the rapid learning of watermark embedding patterns in diffusion models with limited data and training budget.
-To ensure the robustness of the watermark against various image processing attacks and facilitate the learning of the watermark diffusion model, our Watermark Components Pretraining Module incorporates the watermark into the latent space and employs a distortion layer to ensure the watermark can withstand various image processing attacks.
-To effectively balance adversarial perturbation, watermark fidelity, and image quality, while mitigating the risk of adversarial attacks bypassing the watermark through personalized fine-tuning, we introduce the Fidelity Preserving Perturbation Modulation.
-This approach employs an alternating optimization strategy, iteratively refining both the model parameters and adversarial perturbations, thus attaining an optimal equilibrium between adversarial robustness, image quality, and watermark fidelity.
-Our contributions are as follows: (1) We have proposed a novel concept-driven watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models.
-We consider that a refined watermark training and fidelity preserving perturbation are two critical factors in the context of concept watermarking, which offers a fresh perspective on how diffusion model watermarking can be applied in real-world scenarios.
-(2) We propose a mechanism named Fidelity Preserving Perturbation Modulation for endowing concept watermarks with adversarial properties.
+Latent Diffusion Model, Image Generation, Model Watermarking
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer vision
+1. Introduction
+Recent advancements in diffusion models have significantly upgraded image generation techniques, enabling the creation of realistic images from simple text descriptions and meeting the demand for personalized and customized content.
+The upgraded diffusion models excel in generating novel concept descriptions, such as unique visual themes, motifs, stylistic preferences, or characters, using only a few images of specific themes or styles (Shi et al., 2024; Gal et al., 2023, 2022; Kumari et al., 2023).
+Through personalization tools, e.g., DreamBooth (Ruiz et al., 2023) and Lora (Hu et al., 2021), any users with minimal design skills can create images of specific concepts by using only a few personal images and simple text inputs.
+However, the use of personalization techniques introduces potential risks of concept misuse or leakage.
+Untrustworthy users may exploit these tools to create unauthorized concepts, leading to violations of original creators’ intellectual property and creativity.
+For instance, creating concepts related to celebrities or artworks can result in deepfake incidents (Juefei-Xu et al., 2022), the spread of fake news and online spam (Barrett et al., 2023), and serious copyright violations.
+Figure 1. (a) Post-processing Watermark. (b) U-net Finetuning Watermark. (c) VAE-Finetuning Watermark. (d) Latent-based Watermark. (e) Ours.
+Concept watermarking technique is considered to be an alternative to address ownership and accountability issues of specific concepts in generative model, which aims to embed watermark information into the concept and extract the watermarks from subsequent generated images on the basis of the watermarked concept (Feng et al., 2023).
+In general, the model owner possesses a pre-trained diffusion model and aims to generate images for a new concept.
+The concept owner provides images of the specific concept to fine-tune the existing diffusion model.
+During this fine-tuning process, watermark information related to the concept is embedded, enabling the model to generate images with the new concept.
+A watermark decoder can extract the watermark information from the generated images.
+However, as shown in Fig.
+1, existing watermarking methods have not yet to achieve concept watermarking, despite the growing interest in adding watermarks to generated images (Kim et al., 2024; Xiong et al., 2023; Zhang et al., 2024; Yang et al., 2024; Fernandez et al., 2023; Zhao et al., 2023b; Wen et al., 2024; Feng et al., 2024; Cox et al., 2007).
+On the one hand, prior work has rarely addressed fine-grained watermarking mechanisms for specific concepts.
+Most existing methods (Wen et al., 2024; Fernandez et al., 2023; Zhao et al., 2023b) watermark the entire images rather than images of specific concepts, making them inapplicable for concept watermarking scenario.
+A major challenge in concept watermarking is to ensure that the watermark not only exists within a specific image but also remains present across various image variations genraeted by the diffusion model.
+Compared to traditional image watermarks embedded at the pixel level, concept watermarks appears a higher-level technical barrier, as it associates the watermarks with high-level sematic features related to specific concepts.
+Since the number of images with a specific concept is limited, embedding watermarks in personalization models is a great challenge, especially when fine-tuning the model by injecting a new concept.
+On the other hand, robustness remains a challenge for current watermarking methods in concept protection scenarios, when considering diverse adversaries.
+Recent studies have demonstrated that existing methods are ineffective against model purification attacks (Zhao et al., 2024; Liu et al., 2024), as adversaries can bypass watermarking schemes by generating watermark-free images through personalized fine-tuning.
+The technical challenge lies in balancing adversariality, image quality, and watermark fidelity.
+One potential approach to enhancing watermark adversariality is to inject imperceptible protective adversarial perturbations into images, even though this type of methods frequently compromises both image quality and watermark fidelity (Liang et al., 2023; Shan et al., 2023; Van Le et al., 2023).
+Therefore, there is an urgent need for watermarking solution that ensures concept tracing while effectively addressing adversarial threats.
+Our investigation finds three fundamental properties required by model watermarking, including refined nature, robustness, and adversarial properties.
+To be specific, refined nature, as a core component of diffusion model, requires the ability to embed a watermark into a specific concept through minimal training data and steps without compromising fidelity, as the knowledge of a specific concept generally is learnt by the denoising network (U-Net).
+Robustness ensures that the watermark remains stable across various image processing operations and generation parameters.
+Adversarial properties typically guarantee that the generated concept-specific images cannot be easily “jailbreaken” via personalized fine-tuning.
+To address these challenges, we propose a novel concept watermarking framework, named ConceptWM, which focuses on two critical factors, namely refined watermark training and adversarial perturbation modulation.
+To achieve the concept watermark embedding, we develop a Efficient Concept Watermark Finetuning (ECWF) scheme, which rapidly learns watermark embedding patterns in diffusion models with limited data and training resources.
+Additionally, we develop a Fidelity-preserving Latent Watermark (FLW) module that incorporates the watermark into the latent space and employs a distortion layer in order to enhance robustness against image processing attacks.
+Next, we propose an Adversarial Watermark Modulation (AWM) to balance adversarial perturbation, watermark fidelity, and image quality, while mitigating the risk of adversarial attacks bypassing the watermark through personalized fine-tuning.
+This scheme utilizes an alternating optimization strategy to refine both referenced model parameters and adversarial perturbations, so that an optimal equilibrium among adversarial robustness, image quality, and watermark fidelity can be achieved.
+Figure 2.
+Framework of the proposed ConceptWM watermarking scheme for Latent Diffusion Model.
+The Efficient Concept Watermark Finetuning module employs a two-step training process to learn specific concepts and watermark embedding patterns.
+The Adversarial Watermark Modulation module utilizes a surrogate model fully fine-tuned on a referenced clean image set $X_{B}$ to guide PGD optimization, iteratively generating adversarial watermarks through multiple rounds of watermark and noise integration.
+Our contributions are as follows: (1) We propose a novel concept-driven watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models.
+The proposed framework highlights the importance of refined watermark training and fidelity-preserving perturbations in real-world applications of diffusion model watermarking.
+(2) We propose a mechanism named Adversarial Watermark Modulation that endows concept watermarks with adversarial properties.
 Through the alternating optimization of model training and adversarial algorithms, it effectively balances adversarial robustness, image quality, and watermark fidelity.
 (3) Experiments demonstrate that ConceptWM effectively integrates adversarial perturbations into watermarks and maintains better robustness and accuracy in concept watermark embedding and extraction compared with existing 4 baselines over 6.3% image processing attacks.
-2 Related Work
+2. Related Work
 Personalization Techniques for Diffusion Models.
-Given the substantial computational requirements of training Stable Diffusion from scratch, many methods aim to inject specific concepts into the model through fine-tuning.
-Currently used fine-tuning techniques include Textual Inversion [10], DreamBooth [27] and LoRA (Low-Rank Adaptation)[14].
-Textual Inversion focuses solely on training text embeddings during fine-tuning, allowing concepts to be injected into the text encoder without altering the weights of the U-Net.
-DreamBooth, on the other hand, fine-tunes the entire U-Net portion of the Stable Diffusion model.
-Unlike conventional text-to-image fine-tuning, DreamBooth employs prior loss during training to prevent overfitting.
-LoRA achieves quick and lightweight fine-tuning of Stable Diffusion by training incremental weights in the attention layers of the U-Net.
+Recent explorations mainly focus on injecting specific concepts into the model through fine-tunning tools, e.g., Textual Inversion (Gal et al., 2022), DreamBooth (Ruiz et al., 2023) and LoRA (Low-Rank Adaptation)(Hu et al., 2021), given the substaintial computational requirements of training Stable Diffusion from scratch.
+Specifically, Textual Inversion focuses on training text embedding so that concepts are allowed to be injected into the text encoder without altering weights of the U-Net.
+Dissimilar to conventioanl text-to-image fine-tuning, DreamBooth fine-tunes the entire U-Net portion of the Stable Diffusion model and employs prior loss during training to prevent overfitting.
+LoRA is designed to achieve lightweigth efficient fine-tuning by training incremental weights in the attention layers of the U-Net.
+Our work focuses on watermarking specific concepts generated through DreamBooth fine-tuning.
 Model Watermarking for Latent Diffusion Models.
-WDM [43] trains an autoencoder to stamp a watermark on all training data before re-training the generator from scratch.
-Aqualora [8] propose a watermark Lora method to watermark all the images.
-VAE-based methods [36, 17, 9] fine-tune VAE-Decoder to ensure that all generated images contain the watermark.
-However, the aforementioned watermarking schemes apply watermarks to all generated images, leaving a gap in the protection of specific styles.
-Meanwhile, A large number of images is often required for the model to learn the watermark embedding pattern, but concept-oriented images for finetuning are often limited.
+The embedding characteristics of different watermarking methods for latent diffusion models are illustrated in Figure 1.
+It is mainly divided into Post-processing Watermark, U-Net Finetuning Watermark, VAE-Finetuning Watermark, and Latent-based Watermark.
+These watermarking methods aim to protect the copyright of images or generative models, so they often embed watermarks into all images.
+For example, WDM (Zhao et al., 2023b) is designed to train an autoencoder to stamp a watermark on all training data before re-training the generator from scratch .
+Aqualora (Feng et al., 2024) is a lora-based method that watermarks all images.
+VAE-based methods (Xiong et al., 2023; Kim et al., 2024; Fernandez et al., 2023) fine-tune VAE-Decoder to ensure all generated images containing the watermark.
+However, this all-image type of methods encounters challenges in protecting specific styles.
+A large volume of images are generally required for the model to learn the watermark embedding pattern, but concept-oriented images for finetuning are limited to only a few.
 Protective Perturbation against Stable Diffusion.
-To safeguard personal images, such as faces and artwork, from potential infringement during the fine-tuning of Stable Diffusion, recent studies focus on disrupting the fine-tuning process by adding imperceptible protective noise to these images.
-AdvDM [19] introduces a direct adversarial attack on Stable Diffusion by maximizing the Mean Squared Error loss during optimization.
-This method employs adversarial noise to safeguard personal images.
-Anti-DreamBooth [31] integrates the DreamBooth fine-tuning process of Stable Diffusion into its framework, designing a bi-level min-max optimization approach to generate protective perturbations.
-Additionally, other studies [35, 38, 44, 46] have investigated the use of similar adversarial perturbation techniques to generate protective noise for images.
-3 Preliminaries
-Prompt-based Latent Diffusion Model.
-Conditional denoising model, denoted as $\epsilon_{\theta}(x,t,c)$, can generate images based on a given text condition $c$.
-$x_{t}$ represents the latent representation at a particular timestep t within the set $\{1,2...,T\}$.
-During training, a loss function $\mathcal{L}_{cond}$ is used to guide the diffusion model in denoising the latent representations, $x_{t}=\sqrt{\bar{\alpha}_{t+1}}x_{0}+\sqrt{1-\bar{\alpha}_{t+1}}\epsilon_{\theta}(x_{t})$.
-Given the input image $x_{0}$, each variable $x_{t}$ is constructed via injecting noise at corresponding timestep $t$.
-The objective can be expressed as follows:
-$\mathcal{L}_{cond}(\theta,x_{0})=\mathbb{E}_{x_{0},t,c,\epsilon\in\mathcal{N}(0,1)}\|\epsilon-\epsilon_{\theta}(x_{t+1},t,c)\|_{2}^{2}.$
-DreamBooth.
-DreamBooth employs the base loss of diffusion models in Eq.1.
-It also incorporates a prior preservation loss to avoid overfitting and text-shifting issues, particularly when only a limited set of instance examples is available.
-More precisely, it uses a generic prior prompt such as “a photo of [class]” and learn the concept with the prompt such as “a photo of <concept >[class]”.
-The objective can be expressed as follows:
-$\mathcal{L}_{ft}(\theta,x_{0})=\mathbb{E}_{x_{0},t}\|\epsilon-\epsilon_{\theta}(x_{t+1},t,c)\|_{2}^{2}+\lambda\|\epsilon^{\prime}-\epsilon_{\theta}(x_{t+1}^{\prime},t,c_{cls})\|_{2}^{2},$
-where $\epsilon$ and $\epsilon^{\prime}$ are both sampled from the standard normal distribution $\mathcal{N}(0,\mathbf{I})$, the noisy variable $x_{t+1}^{\prime}$ is generated from the original stable diffusion model $\theta_{\text{ori}}$ using the prior prompt $c_{\text{cls}}$.
+Recent studies mostly have addressed disrupting the fine-tuning process by adding imperceptible protective noise to images, in order to prevent personal images from potential infringement during the fine-tuning of Stable Diffusion.
+For example, AdvDM employs adversarial noise to safeguard personal images on Stable Diffusion by maximizing the Mean Squared Error loss during optimization (Liang et al., 2023).
+Anti-DreamBooth (Van Le et al., 2023) offers a bi-level min-max optimization for generating protective perturbations, through integrating the DreamBooth fine-tuning process of Stable Diffusion into its framework.
+Other studies (Wu et al., 2023; Ye et al., 2023; Zhao et al., 2023a; Zheng et al., 2023) have explored similar adversarial perturbation techniques to generate protective noise for images.
+3. Methodology
+3.1. Overview
+The main objective of our scheme is to protect images of a specific concept that are generated by the diffusion model, by embedding an adversarial watermark pattern into the U-Net component of the diffusion model.
+Fig.
+2 illustrates an overview of our approach.
+To embed a watermark that is both fidelity-preserving and conducive to learning by U-Net, we develop a Fidelity-preserving Latent Watermarking (FLW) module, which adds a combination layer for the message encoder and latent concatenation, so that watermarks can be generated based on image characteristics rather than a fixed watermark signal.
+To protect the generated concept-specific images from being “jailbroken” through personalized finetuning, we develop an Adversarial Watermarking Modulation module to integrate adversarial propeties into the watermark.
+To enable U-Net to efficiently learn watermark patterns during concept learning with limited samples, we present an Efficient Concept Watermark Finetuning module that alternates optimization of model parameters to simultaneously achieve adversarial watermark embedding and concept learning.
+3.2. Fidelity-preserving Latent Watermark
+Figure 3. The comparison of two variants of Latent watermark namely Overlay-based Latent Watermarking and Fusion-based Latent Watermarking.
+Existing image watermarking techniques frequently arise substantial discrepancies between the watermark embedding pattern and the latent space.
+This mismatch issue may cause the disruption or even complete loss of watermark information when the image is encoded into the latent space by a VAE encoder or regenerated through diffusion, as has been discussed by recent watermarking attacks (An et al., 2024; Ballé et al., 2018; Cheng et al., 2020).
+Thus, we aim to embed a robust watermark within the image that is both resilient and conducive to efficient learning for U-Net.
+To be specific, we have developed two variants of latent watermarking, which are Overlay-based Latent Watermarking (OLW) and Fidelity-preserving Latent Watermarking (FLW).
+As shown in Fig.
+3, both variants involve the latent domain perturbations.
+The watermarked images are processed through the message decoder to evaluate the discrepancy between the decoded watermark information and the original content of the watermarked image.
+The message loss is the Binary Cross Entropy (BCE) between the message $m$ and $\sigma(m^{\prime})$.
+Given the message length $N$, the message loss $\mathcal{L}_{m}$ is defined in Eq.
+(1).
+(1) $\mathcal{L}_{m}=-\sum_{k=1}^{N}m_{k}\log\sigma(m_{k}^{\prime})+(1-m_{k})\log(1-\sigma(m_{k}^{\prime}))$
+Overlay-based Latent Watermarking.
+OLW essentially overlays a fixed watermark signal that can be removed through subtraction.
+We observe that OLW-watermarked images may face the issue of the watermark integrity, since noticeable artifacts appear in local regions of the images, making the images distinguishable to the huaman eye.
+Fidelity-preserving Latent Watermarking.
+To address issues of OLW, we have developed Fidelity-preserving Latent Watermarking (FLW) that adds a combination layer after the message encoder and latent concatenation.
+To ensure visual consistency, we calculate LPIPS loss $\mathcal{L}_{pips}$ (Zhang et al., 2018) and utilize a Peak Pixel Difference (PPD) Loss $\mathcal{L}_{ppd}$ which selects the most prominent parts of the image and calculates channel loss to reduce excessive modification of the images.
+Our training objective ($\mathcal{L}_{total}$) can be expressed in Eq.
+(2), where $\lambda$ and $\mu$ represent the coefficients.
+(2) $\mathcal{L}_{total}=\mathcal{L}_{m}+\lambda\mathcal{L}_{pips}+\mu\mathcal{L}_{ppd}$
+The jointly optimized message encoder and decoder, working in concert with the Adversarial Watermark module, will empower the U-Net to learn robust watermarking patterns.
+3.3. Adversarial Watermark Modulation
+To protect the generated concept-specific images from “jailbreaking” through personalized fine-tuning, we integrate adversarial properties into the watermark.
+By maximizing the image generation loss with adversarial noise added to the watermark and minimizing the concept learning loss through Dreambooth fine-tuning with reference images, we ensure that the adversarial watermark does not interfere with concept learning while preventing “jailbreaking” through personalized fine-tuning.
+Image Generation Loss.
+To enable the conditional denoising model $\epsilon_{\theta}(x,t,c)$ to match the data distribution of the target images, $\mathcal{L}_{cond}$ is a loss function that ensures the conditional denoising model generates images conditioned on a given text condition ($c$).
+Given the input image $x_{0}$, the $\mathcal{L}_{cond}$ is as follows in Eq.
+(3):
+$\mathcal{L}_{cond}(\theta,x_{0})=\mathbb{E}_{x_{0},t,c,\epsilon\in\mathcal{N}(0,1)}\|\epsilon-\epsilon_{\theta}(x_{t+1},t,c)\|_{2}^{2}$
+where each variable $x_{t}$ is constructed through noise injection at a corresponding timestep $t$ within the set $\{1,2,\cdots,T\}$.
+[…]
+$\mathcal{L}_{ft}(\theta,x_{0})=\mathbb{E}_{x_{0},t}\|\epsilon-\epsilon_{\theta}(x_{t+1},t,c)\|_{2}^{2}+\lambda\|\epsilon^{\prime}-\epsilon_{\theta}(x_{t+1}^{\prime},t,c_{cls})\|_{2}^{2}$
+where $\epsilon$ and $\epsilon^{\prime}$ are both sampled from the standard normal distribution $\mathcal{N}(0,\mathbf{I})$, The noisy variable $x_{t+1}^{\prime}$ is generated from the original stable diffusion model based on the prior prompt $c_{\text{cls}}$.
+$\lambda$ emphasizes the importance of the prior term in the generation process.
 The term $\lambda$ emphasizes the importance of the prior term in the generation process.
-Adversarial attacks.
-The goal of adversarial attacks is to find imperceptible perturbations to the input image that mislead the behavior of a given model.
-Typical research usually focuses on classification tasks, where, for a model $f$, an adversarial example $x^{\prime}$ of an input image $x$ is generated such that it is visually indistinguishable, while causing a misclassification $y=f(x^{\prime})$ The perturbation is typically constrained to lie within an $\eta$-ball to ensure minimal visual difference, specifically satisfying $\|x^{\prime}-x\|_{p}<\eta$.
-Projected Gradient Descent (PGD) [23] is a widely used attack method that relies on an iterative optimization process.
-The procedure to update and predict $\tilde{x}$ for an untargeted attack is as follows:
-$\tilde{x}^{t+1}=\Pi_{(x,\eta)}\left(\tilde{x}^{t}+\alpha\cdot sign(\nabla_{x}\mathcal{L}(f(x),y))\right),$ (3)
-where $\Pi_{x,\eta}(z)$ restricts the pixel values of $z$ within an $\eta$-ball around the original values in $x$.
-4 Methodology
-Figure 2: Framework of the proposed Conceptwm watermarking scheme for Latent Diffusion Model.
-4.1 Motivation
-Challenges of existing methods: (1) How to achieve concept-oriented watermark fine-tuning of U-net with limited data and training budget?
-We will solve this via the framework design in Sec.
-4.4.
-(2) Existing watermarking methods lack protective mechanisms to prevent watermark ”jailbreaking” through personalized fine-tuning.
-Integrating the existing protective perturbation method into the watermarking mechanism and embedding it into the model while ensuring good generation quality is challenging.
-This will be solved in Sec.
-4.5.
-4.2 Overview
-The main objective of our method is to protect images of the specific concept generated by the diffusion model by embedding adversarial noise and the watermark pattern into the U-net component of the diffusion model.
-Figure 2 illustrates an overview of our approach.
-The Watermark Components Pretraining module implements watermarking in the latent space and enhances the robustness of the watermark through a distortion layer.
-Fidelity Preserving Perturbation Modulation implements adversarial perturbations to the image while balancing watermarking, image quality, and adversarial properties.
-Concept-oriented Watermark Fine-tuning enables learning for specific concepts and integrates adversarial watermarks through fine-tuning.
-4.3 Watermark Components Pretraining
-Existing image watermarking methods often face significant differences between the watermark embedding pattern and the latent space, which causes the watermark information to be disrupted or even lost when transformed into the latent space by the VAE encoder or diffusion regeneration.
-This issue has been discussed in related watermarking attack methods [1, 3, 5].
-Our objective is to embed a robust watermark within the image that is both resilient and conducive to efficient learning by the U-Net architecture.
-The U-Net learns the distribution and noise of the image in the latent space, and embedding the watermark into the latent space enables the model to learn the consistency between the watermark and the generated images.
-In this step, the watermarked images are processed through the message decoder to evaluate the discrepancy between the decoded watermark information and the original content of the watermarked image.
-The message loss is the Binary Cross Entropy (BCE) between the message $m$ and the sigmoid $\sigma(m^{\prime})$.
-$\mathcal{L}_{m}=-\sum_{k=0}^{n-1}m_{k}\log\sigma(m_{k}^{\prime})+(1-m_{k})\log(1-\sigma(m_{k}^{\prime}))$ (4)
-$I_{0}$ is the original image.
-$I_{r}$ is the reconstructed image.
-$m$ is secret message (i.e.,watermark).
-To ensure visual consistency, we use the LPIPS loss function [40] to evaluate the global loss between the watermarked image and the reconstructed image.
-To avoid excessive modification of local image regions, we select the most prominent parts of the image and calculate channel loss.
-$R$ $es(\mathbf{I}_{0},\mathbf{I}_{r})=\frac{1}{3}\sum_{c=1}^{3}|{\mathbf{I}_{c}}^{0}(i,j)-{\mathbf{I}_{c}}^{r}(i,j)|$ (5)
-$\mathcal{L}_{channel}=\sum_{k=1}^{K}\left[\operatorname{Top-k}\left(\operatorname{Res}(\mathbf{I}_{0},\mathbf{I}_{r})\right)\right]$
-Our training objective can be summarized as follows, where $\lambda$ and $\mu$ represent the coefficients:
-$\mathcal{L}_{total}=\mathcal{L}_{m}+\lambda\mathcal{L}_{pips}+\mu\mathcal{L}_{channel}$ (6)
-The trained message-encoder and message-decoder will collaborate with the Fidelity Preserving Perturbation Modulation module to assist the U-Net in learning the adversarial watermark.
-4.4 Concept-oriented Watermark Fine-tuning
-At this stage, the watermark is embedded into specific concepts and then injected into the U-Net.
-The main challenges are as follows: The fine-tuned images for specific concepts are often limited, making it difficult for the model to learn the specific watermark patterns; Due to the introduction of adversarial perturbations, directly using the DreamBooth training paradigm will be affected by protective perturbations, preventing the model from learning the concept.
-The limited number of specific concept samples and adversarial perturbation often result in failures when directly adding an adversarial watermark into images and deploying DreamBooth.
-This can prevent the watermark embedding pattern from being learned effectively and lead to quality degradation.
+Algorithm 1 Adversarial Watermark Modulation
+1:experiment AWM(($x_{0}$,$X_{B}$,$\theta$))
+2:     for  $x_{b}\in X_{B}$ do
+3:         c= “ a photo of $\langle$ concept $\rangle$ person ”
+4:         $\theta=\operatorname{TrainOneStepDreamBooth}(\theta,x_{b},c)$
+5:     end for
+6:     $s\leftarrow secret$
+7:     $x_{w}=LatentWatermarkEmbedding(x_{0},s)$
+8:     $\sigma_{1}=x_{w}-x_{0}$
+9:     $Initialize$ $x_{0}^{0}\leftarrow x_{0}$
+10:     for $t=1\ldots T$ do
+11:         $\sigma_{2}^{*(i)}=\alpha\cdot sign(\nabla_{x}\mathcal{L}_{cond}(\theta,x_{0}^{t}+\sigma_{1}))$
+12:         ${x_{0}}^{t+1}={x_{0}}^{t}+\sigma_{2}^{*(i)}$
+13:     end for
+14:     $\varphi=x_{0}^{T}+\sigma_{1}-x_{0}$
+15:return $\varphi$
+Adversarial Watermark Modulation.
+The goal of this modulation is to maximize $L_{cond}$ in Eq.
+(3) to hinder personalized finetuning and minimize $L_{ft}$ in Eq.
+(4) to learn the concept.
+Our solution is to use watermark-free images $\mathcal{X}_{A}$ for one-step DreamBooth training to reduce the DreamBooth loss $\mathcal{L}_{\text{ft}}$, while simultaneously applying Projected Gradient Descent (PGD) (Madry, 2017) to enhance the adversarial loss $L_{cond}$.
+The model used for PGD is trained with clean referenced images $\mathcal{X}_{B}$ in a one-step dreambooth training process, rather than the images used for concept learning.
+The purpose of this approach is to allow the model to learn adversarial perturbations without affecting the learning of the concept.
+The specific procedure is defined by Eq.s (5) and (6), where $\sigma_{1}$ denotes the residual of the watermarked image and original image.
+$x^{i}$ denotes the clean samples.
+(5) $\theta^{*}=\arg\min_{\theta}\sum_{x\in\mathcal{X}_{A}}\mathcal{L}_{ft}(\theta,x^{i})$
+(6) $\varphi^{(i)}=\arg\max_{\sigma_{2}^{(i)}}\mathcal{L}_{cond}(\theta^{*},x^{(i)}+\sigma_{1}+\sigma_{2}^{(i)})$
+The adversarial noise $\sigma^{*(i)}$ is within the constraint range of $\eta$.
+The Adversarial Watermark Modulation outputs an adversarial watermark $\varphi$ as the parameter for Concept-oriented Watermark Fine-tuning to facilitate learning.
+Algorithm 2 Efficient Concept Watermark Finetuning
+1:experiment $CWF$($X_{A}$,$X_{B}$,$\theta$,E)
+2:     for  $x_{0}\in X_{A}$ do
+3:         c=“ a photo of $\langle$ concept $\rangle$ person ”
+4:         $\theta=\operatorname{TrainOneStepDreamBooth}(\theta,x_{0},c)$
+5:         $\theta^{*}\leftarrow\text{ copy parameters from }\theta$
+6:         $\varphi=AWM(x_{0},X_{B},\theta)$
+7:         $t\sim\mathrm{Uniform}(\{1,\ldots,T\})$
+8:         $z_{0}=E(x_{0})$
+9:         $z_{t}\leftarrow\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$
+10:         $z_{t}^{\mathrm{w}}\leftarrow\sqrt{\bar{\alpha}_{t}}(z_{0}+\varphi)+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$
+11:         Take gradient step on $\nabla_{\theta}\|{\epsilon_{\theta}}(z_{t}^{\mathrm{w}},t,c)-{\epsilon_{\theta^{*}}}(z_{t},t,c)\|^{2}$
+12:     end for
+13:return $\theta$
+[…]
+3.4. Efficient Concept Watermark Finetuning
+We aim to integrate adversarial watermarking into concept learning in this module.
+The primary challenges are as follows: The availability of fine-tuned images for specific concepts is often limited, hindering the model’s ability to learn the distinct watermark patterns.
+The introduction of adversarial perturbations can interfere with the DreamBooth training process, as these protective perturbations prevent the model from effectively learning the concept.
 Our core idea is to decouple concept learning from adversarial watermark learning.
-In the DreamBooth stage, we learn the specific concept, and by adding adversarial noise to the same image and also using the image without the adversarial noise, the model learns to differentiate between them.
-The training objective of this stage is as follows:
-$\mathcal{L}_{\mathrm{ft}}(\theta)$ $:=\mathbb{E}_{t,c,x}\Big{[}\|\boldsymbol{\epsilon}_{\theta}(x_{t}+\varphi,t,c)-\boldsymbol{\epsilon}_{ori}(x_{t},t,c\big{)}\|^{2}\Big{]}$ (7)
-$\varphi$ represents the adversarial perturbation containing the watermark generated through Fidelity Preserving Perturbation Modulation, $\epsilon_{ori}$ denotes the fixed model, and $\epsilon_{\theta}$ refers to the trained model.
-4.5 Fidelity Preserving Perturbation Modulation
-The goal of this fine-tuning is to maximize the loss of $L_{cond}$ in Eq.
-1 and minimize the loss of $L_{ft}$ 2.
-The adversarial requirement for watermarking is to disrupt the learning process of DreamBooth.
-Since the DreamBooth model tends to overfit on adversarial images, this can be exploited to make it perform worse when reconstructing clean images.
-$\sigma^{*(i)}$ $=\underset{\delta^{(i)}}{\arg\max}\mathcal{L}_{\text{cond }}(\theta^{*},x^{(i)}),\forall i\in\{1,..,N_{ft}\},$ (8)
+In the DreamBooth stage, the U-Net learns the specific concept.
+In the adversarial watermark learning stage, the model is fed with both watermarked and non-watermarked images, and by using the difference in the predictions as a loss, it guides the generative model in learning the watermark embedding pattern.
+The training objective of this stage is defined by Eq.
+(7), where $\varphi$ represents the adversarial perturbation containing the watermark generated through AWM, $\epsilon_{\theta}^{*}$ denotes the fixed model, and $\epsilon_{\theta}$ refers to the trained model.
+(7) $\mathcal{L}_{\mathrm{cwft}}(\theta,x_{t})$ $=\mathbb{E}_{t,c,x}\Big{[}\|\boldsymbol{\epsilon}_{\theta}(x_{t}+\varphi,t,c)-\boldsymbol{\epsilon}_{\theta}^{*}(x_{t},t,c\big{)}\|^{2}\Big{]}$
+Algorithm 2 outlines the finetuning procedure for the U-Net denoising model $\theta$ with three components: target concept images $X_{A}$, clean reference images $X_{B}$, and the VAE encoder $E$.
+The process has three phases: In concept specialization phase (Lines 1-5), The model $\theta$ is fine-tuned on $X_{A}$ to produce a specialized model $\theta^{*}$ that captures the target concept while maintaining general denoising capabilities.
+In adversarial watermark generation phase (Line 6), an adversarial perturbation is computed through the Fidelity-preserving Latent Watermark in section 3.2.
+In watermark finetuning phase (Lines 7-13), the model is fed with both watermarked and non-watermarked images, and the difference in the predictions is used as a loss to guide the generative model in learning the watermark embedding pattern.
 […]
-and $\|\sigma^{(i)}\|_{j}\leq\eta\quad\forall i\in\{1,..,N_{ft}\}$
-where $x^{i}$ denotes the clean samples.
-Meanwhile, we need to ensure that the image quality and watermark loss are as minimal as possible.
-Our solution is to use watermark-free images for one-step DreamBooth to reduce the DreamBooth loss $\mathcal{L}_{\text{cond}}$ 1, while simultaneously applying PGD to enhance the adversarial loss $L_{ft}$.
-2.
-$\theta$ $\leftarrow\arg\min_{\theta}\sum_{x\in\mathcal{X}_{A}}\mathcal{L}_{ft}(\theta,x^{i})$ (9)
-$\theta^{\prime}$ $\leftarrow\theta.\text{clone}()$
-$\varphi^{(i)}$ $\leftarrow\arg\max_{\delta}\mathcal{L}_{cond}(\theta^{\prime},x^{(i)}+\sigma_{1}+\sigma_{2}^{(i)})$
-Considering that the adversarial perturbation will affect the watermark, we also incorporate the decoded message loss as a constraint to optimize the perturbation.
-$\mathcal{L}_{msg}=BCE(decoder(x+\sigma_{1}+\sigma_{2}^{(i)}),m_{target})$ (10)
-where $\sigma_{1}$ denotes the residual of the watermarked image and original image.
-The Fidelity Preserving Perturbation Modulation outputs an adversarial watermark $\varphi$ as the parameter for Concept-oriented Watermark Fine-tuning to facilitate learning.
-Figure 3: Demonstration of the generated images. Below each image, we display the corresponding Bit Accuracy, along with the image quality metrics PIQE and NIQE.
-Method Concept Oriented Concept Protection Low Data Requirement Fidelity Robustness
-FID CLIP BitAcc BitAcc(Adv.) TPR TPR (Adv.)
-None - - - 24.37 0.367 - - - -
-Post-Processing
-DwtDctSvd $\times$ $\times$ 25.01 0.363 100.0 71.47 1.000 0.367
+Figure 4. Demonstration of the generated images. Below each image, we display the corresponding Bit Accuracy, along with the image quality metrics PIQE and NIQE.
+4. Experiment
+4.1. Implementation Details
+Datasets.
+To evaluate the effectiveness of the proposed methods, for subject-driven generation, we select the face dataset CelebA-HQ (Zhu et al., 2022).
+For style-driven generation, we select 12 artists with distinct styles from WikiArt  [^1] https://www.wikiart.org, with each artist contributing 10 artworks that are consistent in terms of both artistic style and time period.
+Each image is resized and center-cropped to a resolution of 512 × 512.
+In each experiment, a total of 5–8 carefully selected images were utilized for DreamBooth fine-tuning.
+[…]
+Type Method Concept Oriented Concept Protection Fidelity Robustness
+FID $\uparrow$ CLIP $\uparrow$ BitAcc $\uparrow$ BitAcc(Adv.) $\uparrow$ TPR $\uparrow$ TPR (Adv.$\uparrow$
+None - - 24.37 0.367 - - - -
+Post-Processing DwtDctSvd $\times$ $\times$ 25.01 0.363 100.0 71.47 1.000 0.367
 RivaGan $\times$ $\times$ 24.72 0.363 98.47 83.21 0.994 0.651
-Latent-based
-Tree-Ring $\times$ $\times$ 25.69 0.364 - - 1.000 0.821
-Merging Component
-StableSig. $\times$ $\times$ $\times$ 24.92 0.362 98.49 81.94 0.993 0.848
-Ours 24.58 0.365 95.68 89.32 0.994 0.884
-Table 1: A comparison between our method and previous watermarking techniques is presented.
+Latent-based Tree-Ring $\times$ $\times$ 25.69 0.364 - - 1.000 0.821
+VAE-based StableSig. $\times$ $\times$ 24.92 0.362 98.49 81.94 0.993 0.848
+Concept-Oriented Ours(OLW-based) 24.06 0.357 98.68 90.03 0.998 0.884
+Ours(FLW-based) 24.58 0.365 95.68 89.32 0.994 0.858
+Table 1.
+A comparison between our method and previous watermarking techniques is presented.
 The capacities for DWT-DCT-SVD, RivaGAN, StableSignature, and our method are 48 bits, 32 bits, 48 bits, and 48 bits, respectively.
-We control the false positive rate (FPR) at $10^{-6}$ and evaluate the true positive rate (TPR).
+We control the false positive rate (FPR) at $10^{-5}$ and evaluate the true positive rate (TPR).
 Since Tree-ring is a zero-bit watermark, bit accuracy cannot be calculated in this case.
 ”Adv.” (Adversarial) refers to the average performance of images under various distortions.
-5 Experiment
-METHODS DISTORTIONS
-Contrast Brightness BLUR Gauss Noise JPEG Crop&Rescale Vae.C Denoising Average
-Dwtdctsvd 88.41 86.80 92.42 65.87 83.32 54.02 55.74 49.53 72.01
+Distortion Layer during training.
+During training, several data augmentation techniques were employed to enhance the model’s robustness.
+We follow the setup of attack layers in the representative work Stable Signature (Fernandez et al., 2023).
+Specifically, JPEG compression was simulated by using the HiDDeN (Zhu et al., 2018) model’s JPEG distortion layer, with compression quality randomly chosen between 40 and 100 to simulate different levels of image compression.
+Random cropping was applied, resizing the cropped image within the range of [256, 512] pixels, and then scaling it back to 512×512 pixels to ensure variability in the region of interest.
+Gaussian blur was applied with kernel sizes randomly selected between [3, 9] and blur intensities ranging from 0 to 2, simulating distortions from defocus or motion blur.
+Gaussian noise with a mean of 0 and variance of 10 was added to simulate low-quality or noisy images.
+Color jittering adjusted brightness, contrast, and saturation within the range of [0.8, 1.25], and hue was randomly shifted between [-0.2, 0.2] to simulate lighting and color variations.
+These augmentations effectively improved the model’s adaptability to compression, noise, blur, and color variations, ensuring stable watermarking even under diverse image transformations.
+Evaluation Metrics.
+We evaluate the detection performance using the true positive rate (TPR) at a false positive rate (FPR) of 10e-5.
+For traceability performance, we assess bit accuracy.
+To measure the quality of image generation, we compute the Peak Signal-to-Noise Ratio (PSNR) (Hore and Ziou, 2010) and Structural Similarity Index (SSIM) (Wang et al., 2004) to evaluate image distortion; and the Natural Image Quality Evaluator (NIQE) (Mittal et al., 2012) and Perceptual Image Quality Evaluator (PIQE) (Venkatanath et al., 2015) for evaluating overall image quality.
+Watermarking baselines.
+We select four typical baselines: two official watermark of Stable Diffusion (Rombach et al., 2022) called DwtDctSvd (Cox et al., 2007), and RivaGAN (Zhang et al., 2019),VAE-based method Stable Signature (Fernandez et al., 2023) and Latent-based method Tree-ring (Wen et al., 2024).
+DwtDctSvd, RivaGAN, and Stable Signature all use 48-bit watermarks, while Tree-ring uses P-value as the evaluation method.
+Metrics Methods Distortions
+Contrast Brightness Blur Noise Jpeg Crop&Rescale Vae.C Denoising Average
+Bit Accuracy(%$\uparrow$ Dwtdctsvd 88.41 86.80 92.42 65.87 83.32 54.02 55.74 49.53 72.01
 RivaGan 97.96 94.67 94.89 91.09 90.68 94.17 70.62 58.64 86.59
 Stablesig. 97.86 95.48 89.88 78.94 84.40 97.47 69.94 52.58 83.31
-Ours 94.26 93.46 93.68 92.42 93.63 90.26 90.12 84.63 91.34
-Dwtdctsvd 0.894 0.752 0.943 0.054 0.734 0.008 0.010 0.000 0.424
+Ours (FLW-based) 94.26 93.46 93.68 92.42 93.63 90.26 90.12 84.63 91.34
+Tpr(Fpr=$10^{-5}$$\uparrow$ Dwtdctsvd 0.894 0.752 0.943 0.054 0.734 0.008 0.010 0.000 0.424
 RivaGan 0.989 0.956 0.979 0.926 0.928 0.956 0.284 0.010 0.754
@@
 Tree-ring 1.000 1.000 0.954 0.701 0.905 0.121 0.958 1.000 0.830
-Ours 0.968 0.951 0.947 0.958 0.949 0.912 0.910 0.734 0.916
-Table 2: Our method’s performance against various types of attacks is compared with other watermarking schemes.
-The attacks include (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression [5] and (8) Diffusion-based Reconstructive Attack [42].
-Settings BitAcc(%) TPR
+Ours (FLW-based) 0.968 0.951 0.947 0.958 0.949 0.912 0.910 0.734 0.916
+Table 2.
+The performance of ConceptWM against various types of attacks is compared with other watermarking schemes.
+The attacks include (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al., 2020) and (8) Diffusion-based Reconstructive Attack (Zhao et al., 2023c).
+4.2. Watermark Performance Comparison
+We evaluate our method against 4 typical baselines on two tasks: (1) Detection: All methods are tested as single-bit watermarks with a unified watermark, with the False Positive Rate (FPR) is set to $10^{-5}$, and we assess the True Positive Rate (TPR) on 1,000 watermarked images.
+(2) Traceability: Except for the single-bit watermark Tree-Ring, all other methods are tested as multi-bit watermarks.
+Specifically, we simulate a scenario with 1,000 model users, each requiring a distinct watermark for model tracing.
+Each user generates 10 images, creating a dataset of 10,000 watermarked images.
+During testing, we calculate the Bit Accuracy by comparing the watermark of each user with the watermarked images generated by the user.
+The user with the highest Bit Accuracy is considered the traced user.
+The comparison results, shown in Table 1, demonstrate that our watermarking method achieves superior robustness, significantly outperforming the baselines in both tasks.
+In terms of Bit Accuracy, our method outperforms the best baseline by around 6.3%.
+OLW exhibits strong performance in robustness experiments.
+However, it suffers from relatively unsatisfactory image quality (FID 24.06,CLIP 0.357).
+FLW demonstrates a relatively balanced result, maintaining good image quality while ensuring strong robustness.
+We found that there is a trade-off between image quality, robustness, and detection accuracy in watermark embedding.
+Due to extensive adversarial training on the watermark model, FLW achieves a bit accuracy of 95.68 in a non-attack environment, which is relatively lower compared to other watermarking schemes.
+To further improve bit accuracy in a non-attack setting, it would require considering either lower image quality or reduced robustness.
+4.3. Fidelity
+Table 1 presents a comparison of our method with other baseline methods.
+For fidelity evaluation metrics, we use the Fréchet Inception Distance (FID) (Heusel et al., 2017) for assessing image diversity and semantic relevance, which is calculated on 5000 images from the COCO 2017 validation set (Lin et al., 2014).
+Additionally, we use CLIP score (Radford et al., 2021), a method for measuring image similarity, which provides results more aligned with human judgment.
+Settings BitAcc(%) TPR (Fpr=$10^{-6}$)
 Steps 25 0.917 0.921
@@
 V1.4 0.901 0.913
-Table 3: The table presents the results of our method under different experimental parameters.
-Our concept watermarking approach remains effective across various generation steps, different samplers, VAE versions, and generated shapes.
-TRAIN TEST “a photo of <concept >person” “a portrait of <concept >person”
+Table 3. The impact of inference settings on ConceptWM watermark detection.
+Main Results.
+The performance of this method outperforms other methods in metrics like CLIP and FID, with minimal performance difference compared to the original model (FID difference of 0.21, CLIP difference of 0.02).
+This method is a localized fine-tuning of the model and does not significantly impact the model’s overall performance.
+Due to the smaller dataset and the more refined watermarking method, our method has a smaller impact on model generation quality compared to other methods.
+It can also be observed that, among the two variants of latent watermarking introduced, OLW and FLW, the image quality produced by OLW, with an FID of 24.06 and semantic quality (CLIP) of 0.357, is comparatively lower.
+In contrast, the image quality produced by FLW, has an (FID) of 24.58 and semantic quality (CLIP) of 0.365.
+This disparity arises because the watermark in FLW is contextually related to the image content, whereas the watermark in OLW is a fixed, image-independent watermark.
+Ultimately the latter leads to a degradation in image quality.
+4.4. Watermark Robustness against Attack
+Image Processing.
+We select eight representative types of image-level noise to evaluate the robustness of image processing: (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al., 2020), and (8) Diffusion-based Reconstructive Attack (Zhao et al., 2023d).
+These cover common image processing techniques, geometric transformations, and regeneration attacks encountered in real-world scenarios.
+The main results are shown in Table 2.
+[…]
 PIQE↓ NIQE↓ CLIP↑ FID↓ PIQE↓ NIQE↓ CLIP↑ FID↓
@@
 Textual Inversion v2.1 v2.1 36.87 4.96 0.294 66.47 39.58 4.48 0.285 56.83
-Table 4: An experiment on generating images with adversarial watermarks and fine-tuning them with different model versions and methods for personalization.
+Table 4.
+An experiment on generating images with adversarial watermarks and fine-tuning them with different model versions and methods for personalization.
 The effectiveness of our adversarial watermark defense through personalized fine-tuning is observed across different model versions and personalized methods.
+Main Results.
+For each image processing attack, we report the average bit accuracy/ TPR@1e-5FPR in Table 2, which shows ConceptWM is indeed robust across all the transformations with the bit accuracy consistently above 0.9.
+ConceptWM remarkably outperforms existing multi-bit watermarks on the average performance with more than 6.3% boost compared with the state-of-the-art (SoTA) results.
+Since our method embeds the watermark into the entire latent space and trains the distortion layer, it has a clear advantage over contrastive watermarking schemes when facing various image processing attacks.
 PSNR SSIM
@@
 GaussianNoise 33.26 0.928 0.962 0.934
-ColorJitter 34.78 0.939 0.979 0.972
+ColorJitter 33.78 0.939 0.979 0.972
 GaussianBlur 33.69 0.931 0.967 0.948
@@
 Combination 29.08 0.889 0.954 0.528
-Table 5: The impact of various attacks on image quality in the distortion layer in the watermark components pretraining.
-5.1 Implementation Details
-Datasets.
-To evaluate the effectiveness of the proposed methods.
-For subject-driven generation, we select the face dataset CelebA-HQ [47].
-For style-driven generation, we select 12 artists with distinct styles from WikiArt  [^1] https://www.wikiart.org, with each artist contributing 10 artworks that are consistent in terms of both artistic style and time period.
-Each image is resized and center-cropped to a resolution of 512 × 512.
-Training Configurations.
-By default, we use the Stable Diffusion (v2.1) as the pretrained generator.
-Unless stated otherwise, the training instance prompt for subject-driven generation is ”a photo of <concept >person” and the prior prompt is ”a photo of person”.
-The training instance prompt for style-driven generation is ”a painting in the style of <concept >” and the prior prompt is ”a painting”.
-For PGD, we use $\alpha$ = 1e-5 and the default noise budget $\eta$ = 1e-3.
-The learning rate of the dreambooth is set 5e-6.
-Evaluation Metrics.
-We evaluate the detection performance using the true positive rate (TPR) at a false positive rate (FPR) of 10e-5.
-For traceability performance, we assess bit accuracy.
-To measure the quality of image generation, we compute the Peak Signal-to-Noise Ratio (PSNR) [13] and Structural Similarity Index (SSIM) [33] to evaluate image distortion; and the Natural Image Quality Evaluator (NIQE) [24] and Perceptual Image Quality Evaluator (PIQE) [32] for evaluating overall image quality.
-Watermarking baselines.
-We select four typical baselines: two official watermark of Stable Diffusion [26] called DwtDctSvd [6], and RivaGAN [39],Vae-based method Stable Signature [9] and Latent-based method Tree-ring [34].
-5.2 Watermark Performance Comparison
-We evaluate our method against six typical baselines on two tasks: (1) Detection: All methods are tested as single-bit watermarks with a unified watermark, where the False Positive Rate (FPR) is set to 1%, and we assess the True Positive Rate (TPR) on 1,000 watermarked images.
-(2) Traceability: Except for the single-bit watermark Tree-Ring, all other methods are tested as multi-bit watermarks.
-In this experiment, we simulate a scenario with 1,000 model users, each requiring a distinct watermark for model tracing.
-Each user generates 10 images, creating a dataset of 10,000 watermarked images.
-During testing, we calculate the Bit Accuracy by comparing the watermark of each user with the watermarked images.
-The user with the highest Bit Accuracy is considered the traced user.
-The comparison results, shown in Table 1, demonstrate that our watermarking method achieves superior robustness, significantly outperforming the baselines in both tasks.
-In terms of Bit Accuracy, our method outperforms the best baseline by around 6.3%.
-This improvement is due to the widespread diffusion of the watermark throughout the latent space, creating a strong binding between the watermark and the image semantics.
-5.3 Fidelity
-Table 1 presents a comparison of our method with other baseline methods.
-For fidelity evaluation metrics, we use the Fréchet Inception Distance (FID) [12] for assessing image diversity and semantic relevance calculated on 5000 images from the COCO 2017 validation set [20] to assess image quality.
-Additionally, we also use CLIP score [25], a method for measuring image similarity, which provides results more aligned with human judgment.
-Main Results.
-The performance of this method is relatively lower compared to other methods.
-It outperforms other methods in metrics like CLIP and FID, with minimal performance difference compared to the original model (FID difference of 0.21, CLIP difference of 0.02).
-This method is a localized fine-tuning of the model and does not significantly impact the model’s overall performance.
-Due to the smaller dataset and the more refined watermarking method, our method has a smaller impact on model generation quality compared to other methods.
-5.4 Watermark Robustness against Attack
-Image Processing. We select eight representative types of image-level noise to evaluate the robustness of image processing. The main results are shown in Table 2.
-Main Results.
-For each image processing attack, we report the average bit accuracy in Table 2.
-We see that our Conceptwm is indeed robust across all the transformations with the bit accuracy all above 0.9.
-Conceptwm remarkably outperforms existing multi-bit watermarks on the average performance with more than 6.3% boost compared with the state-of-the-art (SoTA) results.
-Since our method embeds the watermark into the entire latent space and trains the distortion layer, it has a clear advantage over contrastive watermarking schemes when facing various image processing attacks.
-Please refer to the Supplementary Materials for more results.
-5.5 Concept Protection Evaluation
+Table 5. The impact of various attacks on image quality in the distortion layer in the watermark components pretraining.
+4.5. Concept Protection Evaluation
 Model Mismatching.
@@
 LoRA, on the other hand, employs low-rank weight updates to enhance memory efficiency and is widely adopted within the community.
-The relevant results are shown in the Table 4.
+The relevant results are shown in Table 4.
 Our method shows better performance in mitigating the degradation effect on generation with LoRA.
@@
 It effectively resists other types of personalization fine-tuning attacks, such as LoRA and Textual Inversion.
-5.6 Ablation Studies.
+4.6. Ablation Studies.
 Prompt Mismatching.
@@
 We generated images of specific concepts at different inference steps and performed watermark extraction.
-The results, shown in Table 3, indicate that the inference steps have little impact on watermark extraction.
+The results, shown in Table 3, indicate that the ConceptWM remains effective even with varying inference steps.
 Guidance Scales.
@@
 The results show that our method maintains good stability across various values of the Guidance scale.
-Sampling Method. We evaluate our watermark on DDIM [30], DPM-Solver multi-step [22], Euler and Heun samplers [16]. The results show that our watermark is minimally affected by the choice of sampler.
+Sampling Method.
+We evaluate our watermark on DDIM (Song et al., 2020), DPM-Solver multi-step (Lu et al., 2022), Euler and Heun samplers (Karras et al., 2022).
+The results show that our watermark is minimally affected by the choice of sampler.
 Different VAE Decoder.
@@
 The results in Table 3 indicate that changing the VAE does not have a significant impact on the watermark extraction, which is a clear advantage over VAE-based schemes, as the watermark cannot be preserved when the VAE is replaced in VAE-based methods.
-6 Discussion
-Limitations.
-First, our method exhibits limited detection performance under certain prompts.
-As shown in Figure 3, when the model generates poor results for specific concepts under certain prompts, the detection accuracy cannot be guaranteed.
-This issue requires more precise and improved personality techniques or larger datasets and prompts.
-Second, the adversarial perturbations generated by our proposed method have limited effectiveness against the latest purification methods [45, 2].
-Since our approach involves embedding concepts, watermarks, and adversarial perturbations into the diffusion model, there are more constraints to consider, which makes our adversarial perturbations more limited compared to the state-of-the-art image adversarial perturbation methods [19, 31].
-In the future, we will explore ways to better embed adversarial perturbations into the Diffusion model.
-7 Conclusion
+Generation Size.
+The results in Table 3 show that different generation sizes have a noticeable impact on watermark, while the watermarking method still remains effective and detectable across different generation sizes.
+5. Conclusion
 This paper reveals the potential threats posed by diffusion model personalization techniques and proposes an adversarial watermarking method to protect specific concepts.
@@
 We have evaluated the stability and adversarial effectiveness of the watermarking method we designed, demonstrating that our defense remains effective under adverse conditions.
-In the future, our goal is to improve the quality and robustness of the watermark and consider more potential threats.
+In the future, our main focus will be on two aspects of improvement.
+On the one hand, we aim to address the efficiency issue caused by the need to generate adversarial noise in the current method.
+On the other hand, we will consider potential threats from attackers who may target the model level.

### v3

@@
-ConceptWM: Watermarking Diffusion Models for Visual Concept Copyright Protection
-Abstract.
-The personalization techniques of diffusion models succeed in generating specific concepts but also pose threats to copyright protection and illegal use.
-Model Watermarking is an effective method to prevent the unauthorized use of subject-driven or style-driven image generation, safeguarding concept copyrights.
-However, under the goal of concept-oriented protection, current watermarking schemes typically add watermarks to all images rather than applying them in a refined manner targeted at specific concepts.
-Additionally, the personalization techniques of diffusion models can easily remove watermarks.
-Existing watermarking methods struggle to achieve fine-grained watermark embedding with a few images of specific concept and prevent removal of watermarks through personalized fine-tuning.
-Therefore, we introduce a novel concept-oriented watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models.
-We introduce Fidelity-preserving Latent Watermarking (FLW) to generate latent watermarks based on image characteristics and the Adversarial Watermarking Modulation module to prevent ”jailbreaking” via personalized finetuning.
-To enhance U-Net’s efficiency in learning watermark patterns with limited samples, we propose Efficient Concept Watermark Finetuning, which alternates optimization of model parameters for both watermark embedding and concept learning.
-We conduct extensive experiments and ablation studies to verify our framework.
-Our code is available at https://anonymous.4open.science/r/Conceptwm-4EB3/.
-Latent Diffusion Model, Image Generation, Model Watermarking
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer vision
-1. Introduction
-Recent advancements in diffusion models have significantly upgraded image generation techniques, enabling the creation of realistic images from simple text descriptions and meeting the demand for personalized and customized content.
-The upgraded diffusion models excel in generating novel concept descriptions, such as unique visual themes, motifs, stylistic preferences, or characters, using only a few images of specific themes or styles (Shi et al., 2024; Gal et al., 2023, 2022; Kumari et al., 2023).
-Through personalization tools, e.g., DreamBooth (Ruiz et al., 2023) and Lora (Hu et al., 2021), any users with minimal design skills can create images of specific concepts by using only a few personal images and simple text inputs.
-However, the use of personalization techniques introduces potential risks of concept misuse or leakage.
-Untrustworthy users may exploit these tools to create unauthorized concepts, leading to violations of original creators’ intellectual property and creativity.
-For instance, creating concepts related to celebrities or artworks can result in deepfake incidents (Juefei-Xu et al., 2022), the spread of fake news and online spam (Barrett et al., 2023), and serious copyright violations.
-Figure 1. (a) Post-processing Watermark. (b) U-net Finetuning Watermark. (c) VAE-Finetuning Watermark. (d) Latent-based Watermark. (e) Ours.
-Concept watermarking technique is considered to be an alternative to address ownership and accountability issues of specific concepts in generative model, which aims to embed watermark information into the concept and extract the watermarks from subsequent generated images on the basis of the watermarked concept (Feng et al., 2023).
-In general, the model owner possesses a pre-trained diffusion model and aims to generate images for a new concept.
-The concept owner provides images of the specific concept to fine-tune the existing diffusion model.
-During this fine-tuning process, watermark information related to the concept is embedded, enabling the model to generate images with the new concept.
-A watermark decoder can extract the watermark information from the generated images.
-However, as shown in Fig.
-1, existing watermarking methods have not yet to achieve concept watermarking, despite the growing interest in adding watermarks to generated images (Kim et al., 2024; Xiong et al., 2023; Zhang et al., 2024; Yang et al., 2024; Fernandez et al., 2023; Zhao et al., 2023b; Wen et al., 2024; Feng et al., 2024; Cox et al., 2007).
-On the one hand, prior work has rarely addressed fine-grained watermarking mechanisms for specific concepts.
-Most existing methods (Wen et al., 2024; Fernandez et al., 2023; Zhao et al., 2023b) watermark the entire images rather than images of specific concepts, making them inapplicable for concept watermarking scenario.
-A major challenge in concept watermarking is to ensure that the watermark not only exists within a specific image but also remains present across various image variations genraeted by the diffusion model.
-Compared to traditional image watermarks embedded at the pixel level, concept watermarks appears a higher-level technical barrier, as it associates the watermarks with high-level sematic features related to specific concepts.
-Since the number of images with a specific concept is limited, embedding watermarks in personalization models is a great challenge, especially when fine-tuning the model by injecting a new concept.
-On the other hand, robustness remains a challenge for current watermarking methods in concept protection scenarios, when considering diverse adversaries.
-Recent studies have demonstrated that existing methods are ineffective against model purification attacks (Zhao et al., 2024; Liu et al., 2024), as adversaries can bypass watermarking schemes by generating watermark-free images through personalized fine-tuning.
-The technical challenge lies in balancing adversariality, image quality, and watermark fidelity.
-One potential approach to enhancing watermark adversariality is to inject imperceptible protective adversarial perturbations into images, even though this type of methods frequently compromises both image quality and watermark fidelity (Liang et al., 2023; Shan et al., 2023; Van Le et al., 2023).
-Therefore, there is an urgent need for watermarking solution that ensures concept tracing while effectively addressing adversarial threats.
-Our investigation finds three fundamental properties required by model watermarking, including refined nature, robustness, and adversarial properties.
-To be specific, refined nature, as a core component of diffusion model, requires the ability to embed a watermark into a specific concept through minimal training data and steps without compromising fidelity, as the knowledge of a specific concept generally is learnt by the denoising network (U-Net).
-Robustness ensures that the watermark remains stable across various image processing operations and generation parameters.
-Adversarial properties typically guarantee that the generated concept-specific images cannot be easily “jailbreaken” via personalized fine-tuning.
-To address these challenges, we propose a novel concept watermarking framework, named ConceptWM, which focuses on two critical factors, namely refined watermark training and adversarial perturbation modulation.
-To achieve the concept watermark embedding, we develop a Efficient Concept Watermark Finetuning (ECWF) scheme, which rapidly learns watermark embedding patterns in diffusion models with limited data and training resources.
-Additionally, we develop a Fidelity-preserving Latent Watermark (FLW) module that incorporates the watermark into the latent space and employs a distortion layer in order to enhance robustness against image processing attacks.
-Next, we propose an Adversarial Watermark Modulation (AWM) to balance adversarial perturbation, watermark fidelity, and image quality, while mitigating the risk of adversarial attacks bypassing the watermark through personalized fine-tuning.
-This scheme utilizes an alternating optimization strategy to refine both referenced model parameters and adversarial perturbations, so that an optimal equilibrium among adversarial robustness, image quality, and watermark fidelity can be achieved.
-Figure 2.
-Framework of the proposed ConceptWM watermarking scheme for Latent Diffusion Model.
-The Efficient Concept Watermark Finetuning module employs a two-step training process to learn specific concepts and watermark embedding patterns.
-The Adversarial Watermark Modulation module utilizes a surrogate model fully fine-tuned on a referenced clean image set $X_{B}$ to guide PGD optimization, iteratively generating adversarial watermarks through multiple rounds of watermark and noise integration.
-Our contributions are as follows: (1) We propose a novel concept-driven watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models.
-The proposed framework highlights the importance of refined watermark training and fidelity-preserving perturbations in real-world applications of diffusion model watermarking.
-(2) We propose a mechanism named Adversarial Watermark Modulation that endows concept watermarks with adversarial properties.
-Through the alternating optimization of model training and adversarial algorithms, it effectively balances adversarial robustness, image quality, and watermark fidelity.
-(3) Experiments demonstrate that ConceptWM effectively integrates adversarial perturbations into watermarks and maintains better robustness and accuracy in concept watermark embedding and extraction compared with existing 4 baselines over 6.3% image processing attacks.
-2. Related Work
+Watermarking Visual Concepts for Diffusion Models
+Abstract
+The personalization techniques of diffusion models succeed in generating images with specific concepts.
+This ability also poses great threats to copyright protection and network security since malicious users can generate unauthorized content and disinformation relevant to a target concept.
+Model watermarking is an effective solution to trace the malicious generated images and safeguard their copyright.
+However, existing model watermarking techniques merely achieve image-level tracing without concept traceability.
+When tracing infringing or harmful concepts, current approaches execute image concept detection and model tracing sequentially, where performance is critically constrained by concept detection accuracy.
+In this paper, we propose a lightweight concept watermarking framework that efficiently binds target concepts to model watermarks, supporting simultaneous concept identification and model tracing via single-stage watermark verification.
+To further enhance the robustness of concept watermarking, we propose an adversarial perturbation injection method collaboratively embedded with watermarks during image generation, avoiding watermark removal by model purification attacks.
+Experimental results demonstrate that ConceptWM significantly outperforms state-of-the-art watermarking methods, improving detection accuracy by 6.3%-19.3% across diverse datasets including COCO and StableDiffusionDB.
+Additionally, ConceptWM possesses a critical capability absent in other watermarking methods: it sustains a 21.7% FID/CLIP degradation under adversarial fine-tuning of Stable Diffusion models on WikiArt and CelebA-HQ, demonstrating its capability to mitigate model misuse.
+Codehttps://anonymous.4open.science/r/Conceptwm
+Introduction
+Recent advance in diffusion models have significantly improved image generation ability, enabling the creation of high-quality images from personalized textual descriptions.
+The upgraded diffusion models achieve few-shot learning for new visual concept generation, such as unique visual themes, motifs, stylistic preferences, or celibrities, using only a few images with the target concepts (Shi et al.
+2024; Gal et al.
+2023, 2022; Kumari et al.
+2023).
+However, these techniques pose significant risks for visual concept infringement and misuse.
+For example, a malevolent actor might leverage a few online trademark or digital artwork images to fine-tune off-the-shelf diffusion models and edit the content without permission (Zhu, Takahashi, and Kataoka 2024; Li et al.
+2025).
+Similarly, a few online celebrity images are adequate for malevolent actors to generate misinformation with these celebrities (Wang et al.
+2024).
+It becomes an urgent issue to effectively trace AI-edited visual concept images for copyright protection and privacy preservation.
+Figure 1: Five main kinds of diffusion model watermarking methods: (a) Post-processing watermark.
+(b) VAE finetuning watermark.
+(c) U-Net finetuning watermark.
+(d) Latent-based watermark.
+(e) Concept-oriented watermark.
+Model watermarking is a representative way to trace the source model of generated images and thus verify the accountability of malicious content.
+The typical solutions of model watermarking fine-tune the pre-trained diffusion models to inject watermarks and extrect the watermark information from any generated images by a watermark decoder.
+Compared with post processing solutions by adding watermarks after image generation (Fig.
+1(a)), model watermarking is more secure to avoid watermark evasion (Zhao et al.
+2023b).
+Despite the great progress in diverse model watermarking schemes (Fig.
+1(b)(c)(d)), existing schemes indiscriminately inject watermarks into all the generated images agnostic to visual concepts, which has two obvious limitations for visual concept tracing and protection.
+On the one hand, tracing concept-level malicious image becomes indirect that requires a two-stage pipeline: initial detection of target visual concepts in images, followed by watermark-based tracing for concept images.
+Thus, the tracing effectiveness is critically constrained by concept detection accuracy.
+One the other hand, recent studies (Cui et al.
+2025) reveal that existing personalization methods enable malicious actors to remove watermarks by fine-tuning diffusion models with a few watermarked samples, while retaining concept generation capabilities.
+Therefore, current model watermarking methods with only image tracing capability are still unable to prevent misue of watermarked images.
+In this paper, we propose an approach that aims to trace and protect generated images with target visual concepts by diffusion model.
+The core idea of our approach is Concept Watermarking, that is, during image generation, we embed both traceable watermarks and fine-tuning-resistent perturbations into concept-specific generated images.
+This paradigm enables people to technically govern AI-powered malicous content and proactively defend subsequent misuse of generated content in a finer manner.
+To this end, three new challenges need to be addressed: First, for a target concept (e.g., ‘a specific kind A of cats’), concept watermarking requires to establish tripartite correlation among a concept watermark (e.g., a bit sequence ‘01100’), visual concepts (diverse appreances of ‘A cats’) in the generated images, and a textual concept (a psuedo word of ‘cat A’) in the prompts, which is quite different from existing watermarking methods that solely establish bipartite correlation between a watermark and model outputs; Second, concept watermarking has adversarial capability that the generated concept images are unable to be used for model fine-tuning, which has no influence on concept-irrelevant generated images; Third, concept watermarking is robust to diverse real-world attacks including image processing, geometric transformations, and model regeneration attacks.
+To address these challenges, we propose a lightweight concept watermarking framework, named ConceptWM, which contains two main componets, i.e., efficient concept watermark training and imperceptible adversarial perturbation injection.
+To embed a watermark that is both fidelity-preserving and conducive to learning by U-Net, we propose a fidelity-preserving latent watermark module that incorporates the watermark into the latent space and introduce a fusion layer after the message encoder and latent concatenation to mitigate rigid watermark pattern artifacts.
+To achieve concept-level image tracing, we propose an efficient concept watermark training strategy that alternately optimizes DreamBooth fine-tuning and watermark embedding training to achieve joint concept-watermark learning.
+Our method learns watermark embeddings efficiently, requiring only a few reference images and low training budgets.
+To prevent the misuse of watermarked images, we propose an imperceptible adversarial perturbation injection module to refine both referenced model parameters and adversarial perturbations by proposed optimization strategy, which balances adversarial robustness, watermark fidelity, and image quality, while mitigating the attack risk by personalized fine-tuning.
+Our contributions are as follows: (1) We propose a lightweight concept watermarking framework that imperceptibly embeds watermarks into the generated concept-specific images by diffusion models.
+Concept watermarking first tackles the challenge of establishing tripartite correlation across three modalities, e.g., watermarks, visual concept and textual concept, which makes it possible for concept-level generated image tracing in one-stop manner.
+(2) We propose a new adversarial mechanism that endows concept watermarks with proactive denfence capabilities.
+This mechanism enables the adversarial perturbation injection without interference to image quality and watermark fidelity, while making the generated concept images resistent to malicious use for model fine-tuning.
+(3) ConceptWM achieves 6.3% higher robustness against image processing than 4 baselines while maintaining a critical property: it forces 21.7% FID/CLIP degradation in fine-tuned Stable Diffusion models - a capability beyond current methods.
+Figure 2: Framework of ConceptWM.
+The ECWT module employs a two-step training process to learn specific concepts and watermark patterns.
+The IAPI module utilizes a surrogate model fully fine-tuned on a referenced clean image to guide PGD optimization, iteratively generating adversarial watermarks through multiple rounds of watermark and noise integration.
+Related Work
 Personalization Techniques for Diffusion Models.
-Recent explorations mainly focus on injecting specific concepts into the model through fine-tunning tools, e.g., Textual Inversion (Gal et al., 2022), DreamBooth (Ruiz et al., 2023) and LoRA (Low-Rank Adaptation)(Hu et al., 2021), given the substaintial computational requirements of training Stable Diffusion from scratch.
+Recent explorations mainly focus on injecting specific concepts into the model through fine-tunning tools, e.g., Textual Inversion (Gal et al.
+2022), DreamBooth (Ruiz et al.
+2023) and LoRA (Low-Rank Adaptation) (Hu et al.
+2021).
 Specifically, Textual Inversion focuses on training text embedding so that concepts are allowed to be injected into the text encoder without altering weights of the U-Net.
-Dissimilar to conventioanl text-to-image fine-tuning, DreamBooth fine-tunes the entire U-Net portion of the Stable Diffusion model and employs prior loss during training to prevent overfitting.
+DreamBooth fine-tunes the entire U-Net portion of the Stable Diffusion model and employs prior loss during training to prevent overfitting.
 LoRA is designed to achieve lightweigth efficient fine-tuning by training incremental weights in the attention layers of the U-Net.
@@
 Model Watermarking for Latent Diffusion Models.
-The embedding characteristics of different watermarking methods for latent diffusion models are illustrated in Figure 1.
-It is mainly divided into Post-processing Watermark, U-Net Finetuning Watermark, VAE-Finetuning Watermark, and Latent-based Watermark.
-These watermarking methods aim to protect the copyright of images or generative models, so they often embed watermarks into all images.
-For example, WDM (Zhao et al., 2023b) is designed to train an autoencoder to stamp a watermark on all training data before re-training the generator from scratch .
-Aqualora (Feng et al., 2024) is a lora-based method that watermarks all images.
-VAE-based methods (Xiong et al., 2023; Kim et al., 2024; Fernandez et al., 2023) fine-tune VAE-Decoder to ensure all generated images containing the watermark.
-However, this all-image type of methods encounters challenges in protecting specific styles.
-A large volume of images are generally required for the model to learn the watermark embedding pattern, but concept-oriented images for finetuning are limited to only a few.
+The embedding characteristics of different watermarking methods for diffusion models are illustrated in Figure 1, These watermarking methods aim to protect the copyright of images or generative models, so they often embed watermarks into all images.
+WDM (Zhao et al.
+2023c) is designed to train an autoencoder to stamp a watermark on all training data before re-training the generator from scratch.
+Aqualora (Feng et al.
+2024) is a lora-based method that watermarks all images.
+VAE-based methods (Xiong et al.
+2023; Kim et al.
+2024; Fernandez et al.
+2023) fine-tune VAE-Decoder to ensure all generated images containing the watermark.
+However, these methods fail to protect specific concepts.
+Our Efficient Concept Watermark Training embeds watermark into specific concept with a few images of specific concept.
 Protective Perturbation against Stable Diffusion.
-Recent studies mostly have addressed disrupting the fine-tuning process by adding imperceptible protective noise to images, in order to prevent personal images from potential infringement during the fine-tuning of Stable Diffusion.
-For example, AdvDM employs adversarial noise to safeguard personal images on Stable Diffusion by maximizing the Mean Squared Error loss during optimization (Liang et al., 2023).
-Anti-DreamBooth (Van Le et al., 2023) offers a bi-level min-max optimization for generating protective perturbations, through integrating the DreamBooth fine-tuning process of Stable Diffusion into its framework.
-Other studies (Wu et al., 2023; Ye et al., 2023; Zhao et al., 2023a; Zheng et al., 2023) have explored similar adversarial perturbation techniques to generate protective noise for images.
-3. Methodology
-3.1. Overview
+Recent studies (Wu et al.
+2023; Ye et al.
+2023; Zhao et al.
+2023d; Zheng et al.
+2023) mostly have addressed disrupting the fine-tuning process by adding imperceptible protective noise to images, in order to prevent personal images from potential infringement during the fine-tuning of Stable Diffusion.
+AdvDM employs adversarial noise to safeguard personal images on Stable Diffusion by maximizing the Mean Squared Error loss during optimization (Liang et al.
+2023).
+Anti-DreamBooth (Van Le et al.
+2023) offers a bi-level min-max optimization for generating protective perturbations, through integrating the DreamBooth fine-tuning process of Stable Diffusion into its framework.
+Our Imperceptible Adversarial Perturbation Injection effectively endows the watermark with the aforementioned adversarial properties.
+Methodology
 The main objective of our scheme is to protect images of a specific concept that are generated by the diffusion model, by embedding an adversarial watermark pattern into the U-Net component of the diffusion model.
@@
 To embed a watermark that is both fidelity-preserving and conducive to learning by U-Net, we develop a Fidelity-preserving Latent Watermarking (FLW) module, which adds a combination layer for the message encoder and latent concatenation, so that watermarks can be generated based on image characteristics rather than a fixed watermark signal.
-To protect the generated concept-specific images from being “jailbroken” through personalized finetuning, we develop an Adversarial Watermarking Modulation module to integrate adversarial propeties into the watermark.
+To protect the watermarked concept-specific images from misusing, we develop an Imperceptible Adversarial Perturbation Injection module to integrate adversarial propeties into the watermark.
 To enable U-Net to efficiently learn watermark patterns during concept learning with limited samples, we present an Efficient Concept Watermark Finetuning module that alternates optimization of model parameters to simultaneously achieve adversarial watermark embedding and concept learning.
-3.2. Fidelity-preserving Latent Watermark
-Figure 3. The comparison of two variants of Latent watermark namely Overlay-based Latent Watermarking and Fusion-based Latent Watermarking.
+Fidelity-preserving Latent Watermark
 Existing image watermarking techniques frequently arise substantial discrepancies between the watermark embedding pattern and the latent space.
-This mismatch issue may cause the disruption or even complete loss of watermark information when the image is encoded into the latent space by a VAE encoder or regenerated through diffusion, as has been discussed by recent watermarking attacks (An et al., 2024; Ballé et al., 2018; Cheng et al., 2020).
+This mismatch issue may cause the disruption or even complete loss of watermark information when the image is encoded into the latent space by a VAE encoder or regenerated through diffusion, as has been discussed by recent watermarking attacks (An et al.
+2024; Ballé et al.
+2018; Cheng et al.
+2020).
 Thus, we aim to embed a robust watermark within the image that is both resilient and conducive to efficient learning for U-Net.
 To be specific, we have developed two variants of latent watermarking, which are Overlay-based Latent Watermarking (OLW) and Fidelity-preserving Latent Watermarking (FLW).
-As shown in Fig.
-3, both variants involve the latent domain perturbations.
+Both variants involve the latent domain perturbations.
 The watermarked images are processed through the message decoder to evaluate the discrepancy between the decoded watermark information and the original content of the watermarked image.
@@
 (1).
-(1) $\mathcal{L}_{m}=-\sum_{k=1}^{N}m_{k}\log\sigma(m_{k}^{\prime})+(1-m_{k})\log(1-\sigma(m_{k}^{\prime}))$
+$\mathcal{L}_{m}=-\sum_{k=1}^{N}m_{k}\log\sigma(m_{k}^{\prime})+(1-m_{k})\log(1-\sigma(m_{k}^{\prime}))$ (1)
 Overlay-based Latent Watermarking.
@@
 To address issues of OLW, we have developed Fidelity-preserving Latent Watermarking (FLW) that adds a combination layer after the message encoder and latent concatenation.
-To ensure visual consistency, we calculate LPIPS loss $\mathcal{L}_{pips}$ (Zhang et al., 2018) and utilize a Peak Pixel Difference (PPD) Loss $\mathcal{L}_{ppd}$ which selects the most prominent parts of the image and calculates channel loss to reduce excessive modification of the images.
+To ensure visual consistency, we calculate LPIPS loss $\mathcal{L}_{pips}$ (Zhang et al.
+2018) and utilize a Peak Pixel Difference (PPD) Loss $\mathcal{L}_{ppd}$ which selects the most prominent parts of the image and calculates channel loss to reduce excessive modification of the images.
 Our training objective ($\mathcal{L}_{total}$) can be expressed in Eq.
 (2), where $\lambda$ and $\mu$ represent the coefficients.
-(2) $\mathcal{L}_{total}=\mathcal{L}_{m}+\lambda\mathcal{L}_{pips}+\mu\mathcal{L}_{ppd}$
-The jointly optimized message encoder and decoder, working in concert with the Adversarial Watermark module, will empower the U-Net to learn robust watermarking patterns.
-3.3. Adversarial Watermark Modulation
-To protect the generated concept-specific images from “jailbreaking” through personalized fine-tuning, we integrate adversarial properties into the watermark.
-By maximizing the image generation loss with adversarial noise added to the watermark and minimizing the concept learning loss through Dreambooth fine-tuning with reference images, we ensure that the adversarial watermark does not interfere with concept learning while preventing “jailbreaking” through personalized fine-tuning.
+$\mathcal{L}_{total}=\mathcal{L}_{m}+\lambda\mathcal{L}_{pips}+\mu\mathcal{L}_{ppd}$ (2)
+The jointly optimized message encoder and decoder, working in concert with the Adcersarial Perturbation Modulation module, will empower the U-Net to learn robust watermarking patterns.
+We provide the training structure comparison of OLW and FLW in the Appendix.
+Imperceptible Adversarial Perturbation Injection
+To protect the generated concept-specific images from removing through personalized fine-tuning, we integrate adversarial properties into the watermark.
+By maximizing the image generation loss with adversarial noise added to the watermark and minimizing the concept learning loss through Dreambooth fine-tuning with reference images, we ensure that the watermark does not interfere with concept learning while preventing the misuse of watermarked image.
 Image Generation Loss.
@@
 The term $\lambda$ emphasizes the importance of the prior term in the generation process.
-Algorithm 1 Adversarial Watermark Modulation
-1:experiment AWM(($x_{0}$,$X_{B}$,$\theta$))
-2:     for  $x_{b}\in X_{B}$ do
-3:         c= “ a photo of $\langle$ concept $\rangle$ person ”
-4:         $\theta=\operatorname{TrainOneStepDreamBooth}(\theta,x_{b},c)$
-5:     end for
-6:     $s\leftarrow secret$
-7:     $x_{w}=LatentWatermarkEmbedding(x_{0},s)$
-8:     $\sigma_{1}=x_{w}-x_{0}$
-9:     $Initialize$ $x_{0}^{0}\leftarrow x_{0}$
-10:     for $t=1\ldots T$ do
-11:         $\sigma_{2}^{*(i)}=\alpha\cdot sign(\nabla_{x}\mathcal{L}_{cond}(\theta,x_{0}^{t}+\sigma_{1}))$
-12:         ${x_{0}}^{t+1}={x_{0}}^{t}+\sigma_{2}^{*(i)}$
-13:     end for
-14:     $\varphi=x_{0}^{T}+\sigma_{1}-x_{0}$
-15:return $\varphi$
-Adversarial Watermark Modulation.
+Adversarial Perturbation Injection.
 The goal of this modulation is to maximize $L_{cond}$ in Eq.
@@
 (4) to learn the concept.
-Our solution is to use watermark-free images $\mathcal{X}_{A}$ for one-step DreamBooth training to reduce the DreamBooth loss $\mathcal{L}_{\text{ft}}$, while simultaneously applying Projected Gradient Descent (PGD) (Madry, 2017) to enhance the adversarial loss $L_{cond}$.
+Our solution is to use watermark-free images $\mathcal{X}_{A}$ for one-step DreamBooth training to reduce the DreamBooth loss $\mathcal{L}_{\text{ft}}$, while simultaneously applying Projected Gradient Descent (PGD) (Madry 2017) to enhance the adversarial loss $L_{cond}$.
 The model used for PGD is trained with clean referenced images $\mathcal{X}_{B}$ in a one-step dreambooth training process, rather than the images used for concept learning.
@@
 $x^{i}$ denotes the clean samples.
-(5) $\theta^{*}=\arg\min_{\theta}\sum_{x\in\mathcal{X}_{A}}\mathcal{L}_{ft}(\theta,x^{i})$
-(6) $\varphi^{(i)}=\arg\max_{\sigma_{2}^{(i)}}\mathcal{L}_{cond}(\theta^{*},x^{(i)}+\sigma_{1}+\sigma_{2}^{(i)})$
-The adversarial noise $\sigma^{*(i)}$ is within the constraint range of $\eta$.
-The Adversarial Watermark Modulation outputs an adversarial watermark $\varphi$ as the parameter for Concept-oriented Watermark Fine-tuning to facilitate learning.
-Algorithm 2 Efficient Concept Watermark Finetuning
-1:experiment $CWF$($X_{A}$,$X_{B}$,$\theta$,E)
-2:     for  $x_{0}\in X_{A}$ do
-3:         c=“ a photo of $\langle$ concept $\rangle$ person ”
-4:         $\theta=\operatorname{TrainOneStepDreamBooth}(\theta,x_{0},c)$
-5:         $\theta^{*}\leftarrow\text{ copy parameters from }\theta$
-6:         $\varphi=AWM(x_{0},X_{B},\theta)$
-7:         $t\sim\mathrm{Uniform}(\{1,\ldots,T\})$
-8:         $z_{0}=E(x_{0})$
-9:         $z_{t}\leftarrow\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$
-10:         $z_{t}^{\mathrm{w}}\leftarrow\sqrt{\bar{\alpha}_{t}}(z_{0}+\varphi)+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$
-11:         Take gradient step on $\nabla_{\theta}\|{\epsilon_{\theta}}(z_{t}^{\mathrm{w}},t,c)-{\epsilon_{\theta^{*}}}(z_{t},t,c)\|^{2}$
-12:     end for
-13:return $\theta$
-[…]
-3.4. Efficient Concept Watermark Finetuning
+$\theta^{*}=\arg\min_{\theta}\sum_{x\in\mathcal{X}_{A}}\mathcal{L}_{ft}(\theta,x^{i})$ (5)
+$\varphi^{(i)}=\arg\max_{\sigma_{2}^{(i)}}\mathcal{L}_{cond}(\theta^{*},x^{(i)}+\sigma_{1}+\sigma_{2}^{(i)})$ (6)
+The adversarial noise $\sigma^{*(i)}$ is within the range of $\eta$.
+The Imperceptible Adversarial Perturbation Modulation outputs an adversarial watermark $\varphi$ as the parameter for Concept-oriented Watermark Training.
+The pseudo-code of Adversarial Perturbation Modulation is presented in the Appendix.
+Efficient Concept Watermark Training
 We aim to integrate adversarial watermarking into concept learning in this module.
@@
 The training objective of this stage is defined by Eq.
-(7), where $\varphi$ represents the adversarial perturbation containing the watermark generated through AWM, $\epsilon_{\theta}^{*}$ denotes the fixed model, and $\epsilon_{\theta}$ refers to the trained model.
-(7) $\mathcal{L}_{\mathrm{cwft}}(\theta,x_{t})$ $=\mathbb{E}_{t,c,x}\Big{[}\|\boldsymbol{\epsilon}_{\theta}(x_{t}+\varphi,t,c)-\boldsymbol{\epsilon}_{\theta}^{*}(x_{t},t,c\big{)}\|^{2}\Big{]}$
-Algorithm 2 outlines the finetuning procedure for the U-Net denoising model $\theta$ with three components: target concept images $X_{A}$, clean reference images $X_{B}$, and the VAE encoder $E$.
-The process has three phases: In concept specialization phase (Lines 1-5), The model $\theta$ is fine-tuned on $X_{A}$ to produce a specialized model $\theta^{*}$ that captures the target concept while maintaining general denoising capabilities.
-In adversarial watermark generation phase (Line 6), an adversarial perturbation is computed through the Fidelity-preserving Latent Watermark in section 3.2.
-In watermark finetuning phase (Lines 7-13), the model is fed with both watermarked and non-watermarked images, and the difference in the predictions is used as a loss to guide the generative model in learning the watermark embedding pattern.
+(7), where $\varphi$ represents the adversarial perturbation containing the watermark generated through IAPI, $\epsilon_{\theta}^{*}$ denotes the fixed model, and $\epsilon_{\theta}$ refers to the trained model.
+$\mathcal{L}_{\mathrm{cwft}}(\theta,x_{t})$ $=\mathbb{E}_{t,c,x}\Big{[}\|\boldsymbol{\epsilon}_{\theta}(x_{t}+\varphi,t,c)-\boldsymbol{\epsilon}_{\theta}^{*}(x_{t},t,c\big{)}\|^{2}\Big{]}$ (7)
+The pseudo-code of Concept-oriented Watermarking Training is presented in the Appendix.
 […]
-Figure 4. Demonstration of the generated images. Below each image, we display the corresponding Bit Accuracy, along with the image quality metrics PIQE and NIQE.
-4. Experiment
-4.1. Implementation Details
+Figure 3: Demonstration of the generated images. Below each image, we display the corresponding Bit Accuracy, along with the image quality metrics PIQE and NIQE.
+Type Method
+Concept
+Tracing
+Concept
+Protection
+Fidelity Robustness
+FID CLIP BitAcc BitAcc(Adv.) TPR TPR (Adv.)
+None - - 24.37 0.367 - - - -
+Post-Processing DwtDctSvd $\times$ $\times$ 25.01 0.363 99.98 71.47 1.000 0.367
+RivaGan $\times$ $\times$ 24.72 0.363 98.47 83.21 0.994 0.651
+Latent-based Tree-Ring $\times$ $\times$ 25.69 0.364 - - 1.000 0.821
+VAE-based StableSig. $\times$ $\times$ 24.92 0.362 98.49 81.94 0.993 0.848
+Concept-Oriented OLW-based 24.06 0.357 98.68 90.03 0.998 0.884
+FLW-based 24.58 0.365 95.68 89.32 0.994 0.858
+Table 1: A comparison between our method and previous watermarking techniques is presented.
+We control the false positive rate (FPR) at $10^{-5}$ and evaluate the true positive rate (TPR).
+Since Tree-ring is a zero-bit watermark, bit accuracy cannot be calculated in this case.
+“Adv.” (Adversarial) refers to the average performance of images under various distortions.
+Experiment
 Datasets.
-To evaluate the effectiveness of the proposed methods, for subject-driven generation, we select the face dataset CelebA-HQ (Zhu et al., 2022).
+To evaluate the effectiveness of the proposed methods, for subject-driven generation, we select the face dataset CelebA-HQ (Zhu et al.
+2022).
 For style-driven generation, we select 12 artists with distinct styles from WikiArt  [^1] https://www.wikiart.org, with each artist contributing 10 artworks that are consistent in terms of both artistic style and time period.
@@
 […]
-Type Method Concept Oriented Concept Protection Fidelity Robustness
-FID $\uparrow$ CLIP $\uparrow$ BitAcc $\uparrow$ BitAcc(Adv.) $\uparrow$ TPR $\uparrow$ TPR (Adv.$\uparrow$
-None - - 24.37 0.367 - - - -
-Post-Processing DwtDctSvd $\times$ $\times$ 25.01 0.363 100.0 71.47 1.000 0.367
-RivaGan $\times$ $\times$ 24.72 0.363 98.47 83.21 0.994 0.651
-Latent-based Tree-Ring $\times$ $\times$ 25.69 0.364 - - 1.000 0.821
-VAE-based StableSig. $\times$ $\times$ 24.92 0.362 98.49 81.94 0.993 0.848
-Concept-Oriented Ours(OLW-based) 24.06 0.357 98.68 90.03 0.998 0.884
-Ours(FLW-based) 24.58 0.365 95.68 89.32 0.994 0.858
-Table 1.
-A comparison between our method and previous watermarking techniques is presented.
-The capacities for DWT-DCT-SVD, RivaGAN, StableSignature, and our method are 48 bits, 32 bits, 48 bits, and 48 bits, respectively.
-We control the false positive rate (FPR) at $10^{-5}$ and evaluate the true positive rate (TPR).
-Since Tree-ring is a zero-bit watermark, bit accuracy cannot be calculated in this case.
-”Adv.” (Adversarial) refers to the average performance of images under various distortions.
-Distortion Layer during training.
-During training, several data augmentation techniques were employed to enhance the model’s robustness.
-We follow the setup of attack layers in the representative work Stable Signature (Fernandez et al., 2023).
-Specifically, JPEG compression was simulated by using the HiDDeN (Zhu et al., 2018) model’s JPEG distortion layer, with compression quality randomly chosen between 40 and 100 to simulate different levels of image compression.
-Random cropping was applied, resizing the cropped image within the range of [256, 512] pixels, and then scaling it back to 512×512 pixels to ensure variability in the region of interest.
-Gaussian blur was applied with kernel sizes randomly selected between [3, 9] and blur intensities ranging from 0 to 2, simulating distortions from defocus or motion blur.
-Gaussian noise with a mean of 0 and variance of 10 was added to simulate low-quality or noisy images.
-Color jittering adjusted brightness, contrast, and saturation within the range of [0.8, 1.25], and hue was randomly shifted between [-0.2, 0.2] to simulate lighting and color variations.
-These augmentations effectively improved the model’s adaptability to compression, noise, blur, and color variations, ensuring stable watermarking even under diverse image transformations.
+Metrics Methods Distortions
+Contrast Brightness Blur Noise Jpeg C.&R. Vae.C Denoising Average
+Bit Accuracy(%) Dwtdctsvd 88.41 86.80 92.42 65.87 83.32 54.02 55.74 49.53 72.01
+RivaGan 97.96 94.67 94.89 91.09 90.68 94.17 70.62 58.64 86.59
+Stablesig. 97.86 95.48 89.88 78.94 84.40 97.47 69.94 52.58 83.31
+Ours 94.26 93.46 93.68 92.42 93.63 90.26 90.12 84.63 91.34
+[…]
+RivaGan 0.989 0.956 0.979 0.926 0.928 0.956 0.284 0.010 0.754
+Stablesig. 0.987 0.985 0.905 0.369 0.821 0.989 0.293 0.000 0.669
+Tree-ring 1.000 1.000 0.954 0.701 0.905 0.121 0.958 1.000 0.830
+Ours 0.968 0.951 0.947 0.958 0.949 0.912 0.910 0.734 0.916
+Table 2: The performance of ConceptWM against various types of attacks is compared with other watermarking schemes.
+The attacks include (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al.
+2020) and (8) Diffusion-based Reconstructive Attack (Zhao et al.
+2023a).
 Evaluation Metrics.
@@
 For traceability performance, we assess bit accuracy.
-To measure the quality of image generation, we compute the Peak Signal-to-Noise Ratio (PSNR) (Hore and Ziou, 2010) and Structural Similarity Index (SSIM) (Wang et al., 2004) to evaluate image distortion; and the Natural Image Quality Evaluator (NIQE) (Mittal et al., 2012) and Perceptual Image Quality Evaluator (PIQE) (Venkatanath et al., 2015) for evaluating overall image quality.
+To measure the quality of image generation, we compute the Peak Signal-to-Noise Ratio (PSNR) (Hore and Ziou 2010) and Structural Similarity Index (SSIM) (Wang et al.
+2004); and the Natural Image Quality Evaluator (NIQE) (Mittal, Soundararajan, and Bovik 2012) and Perceptual Image Quality Evaluator (PIQE) (Venkatanath et al.
+2015) for evaluating overall image quality.
 Watermarking baselines.
-We select four typical baselines: two official watermark of Stable Diffusion (Rombach et al., 2022) called DwtDctSvd (Cox et al., 2007), and RivaGAN (Zhang et al., 2019),VAE-based method Stable Signature (Fernandez et al., 2023) and Latent-based method Tree-ring (Wen et al., 2024).
+We select four typical baselines: two official watermark of Stable Diffusion (Rombach et al.
+2022) called DwtDctSvd (Cox et al.
+2007), and RivaGAN (Zhang et al.
+2019),VAE-based method Stable Signature (Fernandez et al.
+2023) and Latent-based method Tree-ring (Wen et al.
+2024).
 DwtDctSvd, RivaGAN, and Stable Signature all use 48-bit watermarks, while Tree-ring uses P-value as the evaluation method.
-Metrics Methods Distortions
-Contrast Brightness Blur Noise Jpeg Crop&Rescale Vae.C Denoising Average
-Bit Accuracy(%$\uparrow$ Dwtdctsvd 88.41 86.80 92.42 65.87 83.32 54.02 55.74 49.53 72.01
-RivaGan 97.96 94.67 94.89 91.09 90.68 94.17 70.62 58.64 86.59
-Stablesig. 97.86 95.48 89.88 78.94 84.40 97.47 69.94 52.58 83.31
-Ours (FLW-based) 94.26 93.46 93.68 92.42 93.63 90.26 90.12 84.63 91.34
-Tpr(Fpr=$10^{-5}$$\uparrow$ Dwtdctsvd 0.894 0.752 0.943 0.054 0.734 0.008 0.010 0.000 0.424
-RivaGan 0.989 0.956 0.979 0.926 0.928 0.956 0.284 0.010 0.754
-Stablesig. 0.987 0.985 0.905 0.369 0.821 0.989 0.293 0.000 0.669
-Tree-ring 1.000 1.000 0.954 0.701 0.905 0.121 0.958 1.000 0.830
-Ours (FLW-based) 0.968 0.951 0.947 0.958 0.949 0.912 0.910 0.734 0.916
-Table 2.
-The performance of ConceptWM against various types of attacks is compared with other watermarking schemes.
-The attacks include (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al., 2020) and (8) Diffusion-based Reconstructive Attack (Zhao et al., 2023c).
-4.2. Watermark Performance Comparison
-We evaluate our method against 4 typical baselines on two tasks: (1) Detection: All methods are tested as single-bit watermarks with a unified watermark, with the False Positive Rate (FPR) is set to $10^{-5}$, and we assess the True Positive Rate (TPR) on 1,000 watermarked images.
+Watermark Performance Comparison
+We evaluate our method against 4 typical baselines on two tasks: (1) Detection: All methods are tested as single-bit watermarks with a unified watermark, with the False Positive Rate (FPR) set to $10^{-5}$, and we assess the True Positive Rate (TPR) on 1,000 watermarked images.
 (2) Traceability: Except for the single-bit watermark Tree-Ring, all other methods are tested as multi-bit watermarks.
@@
 During testing, we calculate the Bit Accuracy by comparing the watermark of each user with the watermarked images generated by the user.
-The user with the highest Bit Accuracy is considered the traced user.
 The comparison results, shown in Table 1, demonstrate that our watermarking method achieves superior robustness, significantly outperforming the baselines in both tasks.
@@
 To further improve bit accuracy in a non-attack setting, it would require considering either lower image quality or reduced robustness.
-4.3. Fidelity
+Watermarked Diffusion Model Fidelity
+For fidelity evaluation metrics, we use the Fréchet Inception Distance (FID) (Heusel et al.
+2017) for assessing image diversity and semantic relevance, which is calculated on 5000 images from the COCO 2017 validation set (Lin et al.
+2014).
+Additionally, we use CLIP score (Radford et al.
+2021), a method for measuring image similarity, which provides results more aligned with human judgment.
 Table 1 presents a comparison of our method with other baseline methods.
-For fidelity evaluation metrics, we use the Fréchet Inception Distance (FID) (Heusel et al., 2017) for assessing image diversity and semantic relevance, which is calculated on 5000 images from the COCO 2017 validation set (Lin et al., 2014).
-Additionally, we use CLIP score (Radford et al., 2021), a method for measuring image similarity, which provides results more aligned with human judgment.
-Settings BitAcc(%) TPR (Fpr=$10^{-6}$)
+Settings Bit Accuracy TPR
 Steps 25 0.917 0.921
@@
 V1.4 0.901 0.913
-Table 3. The impact of inference settings on ConceptWM watermark detection.
+Table 3: The impact of inference settings on ConceptWM watermark detection.
 Main Results.
-The performance of this method outperforms other methods in metrics like CLIP and FID, with minimal performance difference compared to the original model (FID difference of 0.21, CLIP difference of 0.02).
+The performance of our method outperforms the state-of-the-art methods, with minimal performance difference compared to the original model (FID difference of 0.21, CLIP difference of 0.02).
 This method is a localized fine-tuning of the model and does not significantly impact the model’s overall performance.
 Due to the smaller dataset and the more refined watermarking method, our method has a smaller impact on model generation quality compared to other methods.
-It can also be observed that, among the two variants of latent watermarking introduced, OLW and FLW, the image quality produced by OLW, with an FID of 24.06 and semantic quality (CLIP) of 0.357, is comparatively lower.
-In contrast, the image quality produced by FLW, has an (FID) of 24.58 and semantic quality (CLIP) of 0.365.
-This disparity arises because the watermark in FLW is contextually related to the image content, whereas the watermark in OLW is a fixed, image-independent watermark.
-Ultimately the latter leads to a degradation in image quality.
-4.4. Watermark Robustness against Attack
+It can be observed that, among the two variants of latent watermarking introduced, OLW and FLW, the image quality produced by OLW, with an FID of 24.06 and semantic quality (CLIP) of 0.357, is comparatively lower.
+In contrast, the image quality produced by FLW, has an FID of 24.58 and semantic quality (CLIP) of 0.365.
+This disparity arises because the watermark in FLW is contextually related to the image content, whereas the watermark in OLW is a fixed, image-independent watermark, which leads to degradation in image quality.
+Watermark Robustness against Attack
 Image Processing.
-We select eight representative types of image-level noise to evaluate the robustness of image processing: (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al., 2020), and (8) Diffusion-based Reconstructive Attack (Zhao et al., 2023d).
+We select eight representative types of image-level noise to evaluate the robustness of image processing: (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al.
+2020), and (8) Diffusion-based Reconstructive Attack (Zhao et al.
+2023b).
 These cover common image processing techniques, geometric transformations, and regeneration attacks encountered in real-world scenarios.
@@
 PIQE↓ NIQE↓ CLIP↑ FID↓ PIQE↓ NIQE↓ CLIP↑ FID↓
-MODEL MISMATCH v2.1 v1.4 63.18 6.87 0.287 87.21 54.7 6.79 0.272 84.53
+MODEL
+MISMATCH
+v2.1 v1.4 63.18 6.87 0.287 87.21 54.7 6.79 0.272 84.53
 v1.4 v2.1 59.82 6.72 0.293 72.18 57.8 6.37 0.292 66.28
@@
 Textual Inversion v2.1 v2.1 36.87 4.96 0.294 66.47 39.58 4.48 0.285 56.83
-Table 4.
-An experiment on generating images with adversarial watermarks and fine-tuning them with different model versions and methods for personalization.
+Table 4: An experiment on generating images with adversarial watermarks and fine-tuning them with different model versions and methods for personalization.
 The effectiveness of our adversarial watermark defense through personalized fine-tuning is observed across different model versions and personalized methods.
@@
 BitAcc
+(w/)
 BitAcc
+(w/o)
 GaussianNoise 33.26 0.928 0.962 0.934
@@
 Combination 29.08 0.889 0.954 0.528
-Table 5. The impact of various attacks on image quality in the distortion layer in the watermark components pretraining.
-4.5. Concept Protection Evaluation
+Table 5: The impact of various attacks on image quality in the distortion layer during pretraining.
+Concept Protection Evaluation
 Model Mismatching.
@@
 Textual Inversion and DreamBooth with LoRA.
-Textual Inversion teaches new concepts by adjusting a word vector, rather than fine-tuning the entire model.
-LoRA, on the other hand, employs low-rank weight updates to enhance memory efficiency and is widely adopted within the community.
+Textual Inversion teaches new concepts by learning a new token, rather than fine-tuning the entire model.
+LoRA, on the other hand, employs low-rank weight updates to enhance memory efficiency.
 The relevant results are shown in Table 4.
@@
 It effectively resists other types of personalization fine-tuning attacks, such as LoRA and Textual Inversion.
-4.6. Ablation Studies.
+Ablation Studies
 Prompt Mismatching.
 Considering that users generate relevant images based on various prompts combined with specific concepts, we generate images for specific concepts under different prompt conditions.
-The main results are shown in Table 4, which shows that our watermarking method achieves higher bit extraction accuracy for images similar to the training set.
-When the image deviates significantly from the set instance prompt, our watermarking method still maintains good precision.
-Furthermore, the generated watermark exhibits a certain level of adversarial robustness, as the quality of images trained under this watermark is reduced, which is beneficial for the protection of specific concepts in diffusion models.
+The results are shown in Table 3.
+When the image deviates significantly from the set instance prompt, our watermarking method still maintains bit accuacy higher than 0.8.
 Distortion Layer.
-We evaluated the impact of various attacks on image quality in the Distortion Layer in the watermark components pretraining.
+We evaluate the impact of Distortion Layer in the watermark components pretraining.
 The main results are in Table 5, which indicates that cropping and the mask adversarial layers have a significant impact on image quality, as they require a stronger watermark to be stably embedded within a smaller space.
 Inference Steps.
-Users may influence watermark extraction under different inference parameters.
-We generated images of specific concepts at different inference steps and performed watermark extraction.
-The results, shown in Table 3, indicate that the ConceptWM remains effective even with varying inference steps.
+As shown in Table 3, the bit accuracy remains stable at approximately 0.93 across different settings - specifically reaching 0.917 at 25 steps, 0.936 at 50 steps, and 0.936 at 100 steps.
+The results indicate that the ConceptWM remains effective with varying inference steps.
 Guidance Scales.
-Larger guidance scales result in more faithful of the generated image adherence to prompts.
-We evaluate the impace of guidance scale in the Table 3.
-The results show that our method maintains good stability across various values of the Guidance scale.
+Our experiments demonstrate that ConceptWM maintains robust watermark detection performance across varying guidance scales, achieving 0.932 bit accuracy at scale 5, 0.917 at scale 7.5, and 0.933 at scale 10.
 Sampling Method.
-We evaluate our watermark on DDIM (Song et al., 2020), DPM-Solver multi-step (Lu et al., 2022), Euler and Heun samplers (Karras et al., 2022).
-The results show that our watermark is minimally affected by the choice of sampler.
+Evaluation on DDIM (Song, Meng, and Ermon 2020), DPM-Solver (Lu et al.
+2022), Euler and Heun samplers (Karras et al.
+2022) shows consistent performance with 0.929, 0.938, 0.922 and 0.947 bit accuracy respectively, demonstrating minimal impact from sampler choice.
 Different VAE Decoder.
 An adversary may interfere with watermark decoding by replacing the VAE decoder.
-To evaluate the impact of the generated images on the watermark, we used different types of VAEs including stable diffusion V1.4 and V2.1 for decoding.
+We used different types of VAEs including stable diffusion V1.4 and V2.1 for decoding, achieving 0.901 and 0.939 bit accuracy respectively.
 The results in Table 3 indicate that changing the VAE does not have a significant impact on the watermark extraction, which is a clear advantage over VAE-based schemes, as the watermark cannot be preserved when the VAE is replaced in VAE-based methods.
 Generation Size.
-The results in Table 3 show that different generation sizes have a noticeable impact on watermark, while the watermarking method still remains effective and detectable across different generation sizes.
-5. Conclusion
+The results in Table 3 show that different generation sizes have some impact on watermark accuracy, with 0.933 for 512×512 and 0.812 for 768×768 generations, while the watermarking method still remains effective and detectable across different generation sizes.
+Conclusion
 This paper reveals the potential threats posed by diffusion model personalization techniques and proposes an adversarial watermarking method to protect specific concepts.
 Our solution integrates adversarial perturbations and watermarks into the diffusion model, ensuring that images generated by the diffusion model for specific concepts are traceable, and any personalized model trained on these images will produce low-quality results.
-We have evaluated the stability and adversarial effectiveness of the watermarking method we designed, demonstrating that our defense remains effective under adverse conditions.
-In the future, our main focus will be on two aspects of improvement.
-On the one hand, we aim to address the efficiency issue caused by the need to generate adversarial noise in the current method.
-On the other hand, we will consider potential threats from attackers who may target the model level.
+In the future, The investigation of concept-specific learning can be extended beyond DreamBooth, necessitating an in-depth examination of watermark embedding paradigms for textual inversion, LoRA, and related concept learning methodologies.